### v3

# Stability theory over toroidal or Novikov-type base and canonical modifications

Abstract.
We set up a generalization of ubiquitous one-parameter families in algebraic geometry and their use for stability theories ([Mum65, HL14, AHLH23]) to families over toric varieties and their analytic analogues. The language allows us to reformulate degenerations of “irrational” direction in the literature as canonical objects in a unified manner.
Accordingly, we generalize the (semi)stable reduction-type theorem for $\Theta$-stratification in [AHLH23] of Langton type to our higher rank setup. We also establish complex analytic analogue of the results. As an infinitesimal analogue of toric spectrum, we also use Novikov type rings as it gives more canonicity but its use can be avoided logically for readers for readers who prefer not to use such rings.
As applications, we establish the properness part of the moduli of Calabi-Yau cones (cf., also [Od24a]), and also reduce the properness of the moduli of Kähler-Ricci solitons, again to a finite generation type problem in birational geometry.
1. Introduction
In the classical geometric invariant theory [Mum65], the so-called Hilbert-Mumford criterion for stability of an algebro-geometric object, say ${\bf X}$, with respect to an action of a reductive algebraic group $G$, shows that it is enough to analyze degenerations of ${\bf X}$ along actions of $\mathbb{G}_{m}$ which is inside $G$ ($1$-parameter subgroups). The resulting families are isotrivially degenerating families of ${\bf X}$ over $\mathbb{A}^{1}$. In the case when ${\bf X}$ is a polarized projective variety, they are characterized as test configurations by Donaldson [Don02] (cf., also [Don05]).
More precisely, for the Hilbert-Mumford criterion, one associates a GIT weight to each $1$-parameter subgroup and their positivity is equivalent to the stability. Alternatively, it is well known that one can consider maximal tori $T$ of $G$, instead of $1$-parameter subgroups, and encode all GIT weights data within $T$ in the weight polytope. The corresponding natural extension of test configurations or the above type isotrivial degenerations, replacing $\mathbb{G}_{m}$ by a maximal torus, is isotrivial family over $T$ and its partial equivariant compactification i.e., toric varieties. A simple toy model for the principal objects in this paper is an isotrivial degeneration over a higher-dimensional toric variety.
More generally, we introduce more extensive algebraic (partly stack-theoretic) generalities with such perspective, which leads to introduce e.g., isotrivial degenerating families over toroidal bases, Novikov type valuation rings of high rank, and their analytic analogues. In particular, focusing on the context of stability and moduli theories of varieties, we introduce the notion of generalized test configurations and their limit versions. Then, together with a companion paper, we show that the generalized notion is useful to treat the following types of degenerations in “irrational direction” in literatures, a more canonical and unified manners:
  • klt singularities to irregular Ricci-flat cones ([DS17, Li18]),
  • Kähler-Ricci flow on Fano manifolds converging to Kähler-Ricci solitons ([CSW18, HL20]),
  • negative weight deformations to asymptotically conical Calabi-Yau manifolds ([CH22]).
In our §4 and [Od24a, Od24c], we discuss and explore these examples one by one, but until §3 of this paper we focus on the general framework. In particular, our §3 gives some generalization of the $\Theta$-(semi)stability theory and semistable reduction theorem ([HL14] and [AHLH23] respectively), as a core theorem in this paper. Here, recall $\Theta:=[\mathbb{A}^{1}/\mathbb{G}_{m}]$ over a field $k$ is the quotient stack whose coarse moduli is just one point which is used as a source of map to express equivariant isotrivial families over $\mathbb{A}^{1}$. [HL14, AHLH23] introduces the notion of $\Theta$-strata and $\Theta$-stratification on (highly non-separated) algebraic stacks, as a stacky reminicsents of the Hesselink-Kempf-Kirwan-Ness stratification [Kir84, Nes84] and the Harder-Narasimhan filtration of coherent sheaves [HN74]. We call a natural higher $=:r$) version of it as higher $\Theta$-strata by roughly considering degenerations along $\Theta^{r}=[\mathbb{A}^{r}/T]$ with $T:=\mathbb{G}_{m}^{r}$ (temporarily). One can also replace $\mathbb{A}^{r}$ by other affine toric variety $U_{\tau}$ for a cone $\tau$. Also, as an irrational analogue of DVR, we introduce Novikov type spectra $\overline{\Delta_{\xi}}(R)$ (Definition 2.3). This gives an extension of the valuation ring $R$.
To give a feel of our main statements on such higher $\Theta$-stratum, here we show its quick brief outlook, although we refer to later Theorem 3.8, 3.15, Corollary 3.10 for the precise statements and generalizations. Roughly put, it shows the existence of “improvement/modification” of one parameter family, which avoids some “too unstable” locus in a fairly general setup.
For below, we consider a quotient stack $\mathcal{M}=[M/G]$ over $\mathbbm{k}$, where $M$ is an algebraic $\mathbbm{k}$-scheme and $G$ is a linear algebraic $\mathbbm{k}$-group (e.g., ${\rm GL}_{n}$ or algebraic torus). In that case, higher $\Theta$-stratum $\mathcal{Z}^{+}=[Z/G]$ with the closed subscheme $Z\subset M$ and a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ roughly means for each point $p\in Z(\mathbbm{k})$, there is its isotrivial degeneration (the generalized Harder-Narasimhan filtration) along the affine toric variety $U_{\tau}\supset T$ with a homomorphism $T\to G$ and $Z^{+}$ is the maximal locus where these structure varies uniformly. Note that our $r$ is not necessarily the rank of $G$.
Theorem 1.1 (Higher $\Theta$-semistable reduction: cf., Corollary 3.10).
Consider a quotient stack $\mathcal{M}$ over a field $\mathbbm{k}$ and its higher $\Theta$-strata $\mathcal{Z}^{+}\subset\mathcal{M}$ for the cone (see Definition 3.3).
Take any $\mathbbm{k}$-morphism $f\colon\Delta=\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t]]\to\mathcal{M}$ such that its closed point maps into $\mathcal{Z}^{+}\subset\mathcal{M}$ while the generic point maps outside $\mathcal{Z}^{+}$. We denote the restriction of $f$ to $\mathop{\mathrm{Spec}}\nolimits(K)$ as $f^{o}$.
Then, for a finite extension $R^{\prime}(\simeq\mathop{\mathrm{Spec}}\nolimits\mathbbm{k}[[t^{1/d}]]$ for some $d\in\mathbb{Z}_{>0}$) of $R$, there exists a modification of $f$ as follows:
$f_{{\rm new}}\colon\mathop{\mathrm{Spec}}\nolimits(R^{\prime})\to\mathcal{M},$
which still extends $f^{o}$ but now sends the closed point $c_{R^{\prime}}$ to a point outside $Z^{+}$.
If there are finite strata $\{\mathcal{Z}_{i}^{+}\}_{i}$ inside $\mathcal{M}$ which form what we call higher $\Theta$-stratification (Definition 3.17), extending the case #$\#\{i\}=1$ above, we can iterativly apply the above modification (see Theorem 3.18) to show the properness of the complement of $\mathcal{Z}_{i}^{+}$s, which often admit moduli interpretations (see Appendix 4 for specific examples). Indeed, in the moduli contexts, $\mathcal{Z}_{i}$s often appear as unstable locus in the spirit of [Mum65] so that their complement is nothing but the locus which parametrizes semistable objects.
Here are some technical comments in order. For Theorem 1.1 mentioned above, we also refine the statements in the main body of this text from at least the following four different aspects.
The first aspect is about the new base $R^{\prime}$. There are many choices of the pair of DVR $R^{\prime}$ and $f_{\rm new}$ as a family one can take in our construction, though $f_{\rm new}(c_{R^{\prime}})$ is more canonical. Indeed, they roughly corresponds to choice of integral points in a certain rational polyhedral cone so that there is no canonical choice and there are infinite possibilities. Nevertheless, if we fix a possibly irrational vector $\xi(\neq 0)\in\tau$ and allow $R^{\prime}$ to be non-discrete valuation ring in general, there is more canonical choice of $R^{\prime}$ as a bigger valuation ring for each $\xi$ and we expect its importance for future study. That is the Novikov type ring $\overline{\Delta_{\xi}}(R)$ we mentioned above, as we introduce and discuss in §2. We also discuss various variants of related Novikov type rings and relation with toric geometry, motivated by the above reason.
Secondly, we also prove its complex analytic version (Theorem 3.15) by using the universal covering of punctured disk or real ray in the unit disk (cf., Definition 2.4). We call this replacement of $f$ as higher $\Theta$-semistable reduction and call its limit as higher $\Theta$-semistable limit. When $\xi$ is an integral vector, our result partially recovers [AHLH23, §6, Theorem 6.3, Theorem 6.5] by more scheme theoretic approach.
Thirdly, in the case when $\mathcal{M}$ is of the form $[M/T]$, we prove certain uniqueness of the above modification in a weak sense. See (iii) of Theorem 3.8 for the details.  [^1] We hope this will be benefitial for further study of bubbling theory [Od24c].
Lastly, but almost obviously, if there are certain stratifications of $\mathcal{M}$ consists of higher $\Theta$-strata, we can simply repeat the above modification to give a stronger result (Theorem 3.18) as it was in [AHLH23].
The framework and theorems in this paper are applied in our §4 and our companion papers ([Od24a, Od24c]), which are mostly in the context of Kähler geometry and algebraic theory related to K-stability.
Notes added in version 3
About $8$ and a half month after the first version of this paper, we saw some developments i.e., [BHLINK25] appeared, but with a different motivation. At the end of the paper, put in our version 3 on arXiv, we discuss the overlap, interesting relations and certain differences between this paper and [BHLINK25].
2. Generalities on multi-parameter families
This section is of preparatory nature. We generalize some part of the theories of one parameter families and degenerations, which appear in the theory of K-stability or its local analogues for klt singularities (cf., [Li18] and subsequent developments), and the stacky stability theories in [HL14, AHLH23], to those over toroidal base of often higher dimension. Some germs of our theory can be seen in [Spi90, Tei03, DS17, LX18, HL20], for instance. We observe that our language allows re-explanation in a unified manner of those existing K-stability theories.
Firstly, we introduce a notion of generalized test configurations as certain type of multi-parameter isotrivial family over toroidal bases and their analogues, which unifies various families in the field of K-stability and Kähler geometry, such as
  1. ($\mathrm{i}$)
    test configurations and their affine versions ([Don02, CS18, CS19]) cf., Example 2.7,
  2. ($\mathrm{ii}$)
    negative weight deformation in the sense of [CH22, Definition 1.7] cf., Example 2.18 and
  3. ($\mathrm{iii}$)
    each step of the recent $2$-step degeneration theory (originated in [DS17])
    • e.g., the canonical degeneration of klt singularities to its possibly irregular K-semistable Fano cone via the volume-minimizing valuation of klt singularities ([Li18]) in Example 2.9,
    • $\mathbb{R}$-test configurations in [CSW18, HL20] cf., Example 2.14,
    • (the $2$nd steps to the polystable replacements ([DS17, LWX21]) as Examples 2.13),
  4. ($\mathrm{iv}$)
    (or more generally, degeneration of GIT semistable objects to GIT polystable objects ([Mum65]) cf., Remark 2.12).
Further, as we will later explain, the notion allows us to treat the above examples of degenerations ($\mathrm{ii}$) and ($\mathrm{iii}$) in a (more) canonical manner, to remove the broad ambiguity of choices of $1$-parameter subgroups, except for those with parentheses.
2.1. Setup of the bases
Since our bases will be often toric geometric, first we set the usual notation for toric geometry.
Notation 1.
We fix a free abelian group (lattice) $N$ of rank $r$ and its rational polyhedral cone $\tau\subset N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$. The dual lattice of $N$ is denoted as $M:={\rm Hom}(N,\mathbb{Z})$ and we set $M_{\mathbb{R}}:=M\otimes\mathbb{R}$. We define the algebraic split torus $T=\mathbb{G}_{m}\otimes_{\mathbb{Z}}N$ as ${\rm Spec}\mathbb{C}[M]$, where $\mathbb{C}[M]$ means the group ring. If we regard $M$ as the character lattice of $T$, we denote by $\chi(\vec{m})$ the character of $T$ which corresponds to each $\vec{m}\in M$. An affine toric partial compactification of $T$ for $\tau$ is denoted by $U_{\tau}(\supset T)$. We write $\tau^{\vee}:=\{x\in M\otimes\mathbb{R}\mid\langle x,\tau\rangle\subset\mathbb{R}_{\geq 0},$ its interior as $\tau^{\vee,o}$, $\mathcal{S}_{\tau}:=\tau^{\vee}\cap M$, and $\mathcal{S}_{\tau}^{o}:=\mathcal{S}_{\tau}\cap\tau^{\vee,o}$.
We take $(0\neq)\xi\in\tau$ which is not contained in any proper $\mathbb{Q}$-linear subspace of $N_{\mathbb{Q}}:=N\otimes_{\mathbb{Z}}\mathbb{Q}$ i.e., of rational rank $r$. Henceforth, we simply call such $\xi$ as being irrational. The latter is not finitely generated if $r>1$. We introduce the monoids
$M_{\xi\geq 0}$ :=$:=\{m\in M\mid\langle m,\xi\rangle\geq 0\},$
$\Gamma_{\geq 0}$ $:=\langle M_{\geq 0},\xi\rangle\subset\mathbb{R}.$
Here, $\langle$,$\rangle$ , ⟩ means the natural pairing which comes from the definition of $M$. For our later applications to relate with differential geometry, $\xi$ is often rather a given data, rather than $\tau$ (or even $N$).
Also, for any algebraic scheme $V$ over $\mathbb{C}$, $V^{\rm an}$ (or $V(\mathbb{C})^{\rm an}$) refers to the (complex) analytification i.e., $V(\mathbb{C})$ with the complex analytic (euclidean) topology.
We prepare the bases of “irrational deformations” in an algebraic manner, following [Nov81, Nov82, FOOO09, AJ10, AI21] for instance.
Definition 2.1 (Novikov-type rings and spectra).
Under the Notation 1, the global N-ring  [^2] N stands for Novikov as these rings are variants of so-called Novikov ring, in the contexts of Morse theory, Gromov-Witten theory, symplectic geometry, which seem to have different versions in the literatures cf., e.g., [Nov81, Nov82, FOOO09, AJ10, AI21]. over a ring $A$ for $\xi\in N_{\mathbb{R}}$ is the semiring $A[\Gamma_{\geq 0}]$ and we call its spectra $\mathop{\mathrm{Spec}}\nolimits A[\Gamma_{\geq 0}]$ as global N-spectra, which we denote as $\Delta_{\xi}^{\rm gl}(A)$ or simply $\Delta_{\xi}^{\rm gl}$ if the base ring $A$ is obvious from the context. Then, we define the following.
  1. ($\mathrm{i}$)
    Let $I$ be the ideal of $A[\Gamma_{\geq 0}]$ generated by the monomials corresponding to $\Gamma_{\geq 0}$ and the formal completion of $A[\Gamma_{\geq 0}]$ with respect to $I$ is denoted as $A[[\Gamma_{\geq 0}]]$. With its formal variable $t$, the N-ring over a ring $A$ for $\xi\in N_{\mathbb{R}}$ is a subring of $A[[\Gamma_{\geq 0}]]$ defined as
    $A_{\xi}:=\{\sum_{i=1,2,\cdots}a_{i}t^{r_{i}}\mid a_{i}\in A,r_{i}\in\Gamma_{\geq 0},r_{1}<r_{2}<\cdots,r_{i}\to\infty(i\to\infty)\}.$
    We call ${\rm Spec}(A_{\xi})$ the N-spectrum and denote it by $\Delta_{\xi}(A)$. Note that $T$ over ${\rm Spec}(A)$ acts on $\Delta_{\xi}(A)$. Later, we often consider the case when the base $A$ is a field $\mathbbm{k}$ and in that case we denote $A_{\xi}$ as $\mathbbm{k}_{\xi}$. Note that it is clearly a valuation ring of rational rank $r$. When the base field is obvious, we often abbreviate $\Delta_{\xi}(A)$ as $\Delta_{\xi}$.
  2. ($\mathrm{ii}$)
    If $A\subset\mathbb{C}$ (or more generally, if $A$ is associated a valuation) and a semigroup $\Gamma_{\geq 0}$ is a submonoid of $\mathbb{R}_{\geq 0}$, we define the convergent N-ring $A[[\Gamma_{\geq 0}]]^{\rm conv}$ as
    $\{f=\sum_{i=1,2,\cdots}a_{i}t^{r_{i}}\in A_{\xi}\mid\exists\epsilon>0\text{ s.t. }f(t)\text{ absolutely converge for any }t\in(0,\epsilon)\}.$
    We write its fraction field as $A((\Gamma_{\geq 0}))^{\rm conv},$ and denote their spectrum $\mathop{\mathrm{Spec}}\nolimits A[[\Gamma_{\geq 0}]]^{\rm conv}$ (resp., $\mathop{\mathrm{Spec}}\nolimits A((\Gamma_{\geq 0}))^{\rm conv}$) as $\Delta_{\Gamma_{\geq 0}}^{\rm conv}(A)$ (resp., $\Delta_{\Gamma_{\geq 0}}^{*,{\rm conv}}(A)$) and call them convergent N-spectrum. We often consider the case when $\Gamma_{\geq 0}$ is of the form of Notation 1 with some $\xi\in N_{\mathbb{R}}$, and denote them by $\Delta_{\xi}^{\rm conv}(A)$ (resp., $\Delta_{\xi}^{*,{\rm conv}}(A)$).
We can interpret the above notion in terms of the framework of limit toric variety in [Od22b].
Proposition 2.2.
We fix a base integral domain $A$ and irrational $\xi\in N_{\mathbb{R}}$ as Notation 1. Consider the set of all rational polyhedral cones $\tau$ of $N_{\mathbb{R}}$ which includes $\xi$, and associate the obvious partial order with respect to the inclusion.
[…]
Note that each $U_{\tau}(\mathbbm{k})$ for a field $\mathbbm{k}$ has Krull dimension $r$ while the limit $\Delta_{\xi}(\mathbbm{k})$ has dimension $1$, interestingly.
Proof.
Since the projective system $\{U_{\tau}\}_{\tau}$ is that of affine schemes, it is enough to check at the level of rings. Then, it is a straightforward confirmation so we leave the details to the readers. ∎
From the perspective of the above proposition, it is naturally to consider the quotient stack $[\Delta_{\xi}^{\rm gl}/T]$ over the category of (non-finite type) schemes, and denote it as $\Delta_{\xi}$ henceforth. For later purpose in the next section, we also introduce a $R$-algebra variant. More specifically, it will appear as a natural (canonical) base of a modification of morphism from ${\rm Spec}(R)$ to an algebraic stack, in our higher analogue of semistable reduction type theorem (Theorem 3.8, Corollary 3.10).
Definition 2.3 (Novikov-type spectra over $R$).
For $N,\xi$ as notation 1, we fix a local ring $R$ and also fix its non-unit non-zero element $\pi$. We also take and fix $c\in\mathbb{R}_{>0}$, set $\xi^{\prime}:=(\xi,c)\in N_{\mathbb{R}}\times\mathbb{R}_{>0}.$
Consider the minimum lattice $N^{\prime}\subset N\oplus\mathbb{Z}$ whose spanning $\mathbb{R}$-linear subspace contains $\xi^{\prime}$, which we suppose to have rank $r$. This assumption naturally arises in our later application in §3.2. We take the projective system of rational polyhedral cones $\tau^{\prime}$ of $N^{\prime}$ which contains $\xi^{\prime}$. Then consider the natural projection
$p\colon M\oplus\mathbb{Z}\twoheadrightarrow M^{\prime},$
which is dual to $N^{\prime}\hookrightarrow N\oplus\mathbb{Z}$. Then, consider the affine scheme over $R$ as
(1) $U_{\tau^{\prime}}(R):=\mathop{\mathrm{Spec}}\nolimits(R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})),$
and its projective limit
[…]
Morally speaking, it is global in the direction of $\mathbb{G}_{m}\otimes N^{\prime}$-action but it is only local in $R$ direction, which is the reason why we avoid the superscript gl in this case and just use the overline. Here, $\tau^{\prime\vee}$ means the dual cone of $\tau^{\prime}$. As $p(\vec{0},1)\in M^{\prime}\cap\tau^{\prime\vee},$ $\chi^{p(\vec{0},1)}$ means the corresponding monomial (character) of $R[M^{\prime}\cap\tau^{\prime\vee}]$.
We also note that the above ring $R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})$ is a quotient ring of $(R[X])[\tau^{\vee}\cap M]/(X\chi^{p(\vec{0},1)}-\pi)$, where $X$ is an extra variable and the quotient is by the ideal $(X-1)$. We also define
[…]
These and their further extensions will appear in the proofs of semistable reduction type theorems in §3.2 and later. See Definition 3.5, Remark 3.7 Theorem 3.15, and Corollary 3.10.
In the setup we later apply above to Theorem 3.8 and Corollary 3.10, there is a canonical choice of $c$ such that the lift of $\tau$ as $\tau^{\prime}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ contains $(v_{i},1)$ for each of the basis $\{v_{i}\}_{i}$ of the regular generator of the cone $\tau\cap N$ thanks to the irrationality of $\xi$. $\pi$ is also naturally decided in some context later. In such cases, we abbreviate and denote as $\overline{\Delta_{\xi}}(R,\pi)$ (resp., $\widetilde{\Delta_{\xi}}(R,\pi)$), $\overline{\Delta_{\xi}}(R)$ (resp., $\widetilde{\Delta_{\xi}}(R)$), or even $\overline{\Delta_{\xi}}$ (resp., $\widetilde{\Delta_{\xi}}$) for brevity. These are also regarded as versions of N-spectra.
A real ($\mathbb{R}$) analogue of $\Delta_{\xi}$ and $\overline{\Delta}_{\xi}$ is $\tilde{R}_{\xi}$ below.
Definition 2.4 (Analytic analogue).
If the base ring is $A=\mathbbm{k}=\mathbb{C}$, then we consider $r_{\xi}\colon\mathbb{R}\to N_{\mathbb{R}}$ which simply maps $x$ to $x\xi$. $\exp\colon\mathbb{R}\mapsto\mathbb{R}_{>0}\subset\mathbb{C}^{*}=\mathbb{G}_{m}(\mathbb{C})$ means the exponential map $r\mapsto e^{r}$. We also set the notation $(\cdot)^{\xi}:=R_{\xi}\circ\log\colon\mathbb{R}_{>0}\to\mathbb{R}_{>0}\otimes N$, which informally means the $\xi$-th power map. Consider the map
$R_{\xi}:=({\rm id}_{N}\otimes\exp)\circ r_{\xi}\colon\mathbb{R}\to T(\mathbb{C}),$
where ${\rm id}_{N}\colon N\simeq N$ is the identity map. Note that because of the irrationality of $\xi$, this can not be extended to $\mathbb{C}$. We denote the universal cover of $\mathbb{C}^{*}$ as $a\colon\widetilde{\mathbb{C}^{*}}\to\mathbb{C}^{*}$, which is isomorphic to $\mathbb{R}_{>0}\times(\widetilde{U(1)}\simeq\mathbb{R})$, and for each $\epsilon\in\mathbb{R}_{>0}$. ~$\widetilde{\mathbb{C}}_{\epsilon}^{*}$ denotes the preimage of $\{t\in\mathbb{C}^{*}\mid|t|<\epsilon\}$.
Then, $R_{\xi}\circ a$ nevertheless extends to ~$\widetilde{\mathbb{C}^{*}}\to T(\mathbb{C})$ as a holomorphic (analytic) homomorphism between complex Lie groups. We denote it as
$\tilde{R}_{\xi}\colon\widetilde{\mathbb{C}^{*}}\to T(\mathbb{C}).$
Note that the image is not inside the real part $N\otimes\mathbb{R}_{>0}$ as it is also nontrivial along the argument direction $\widetilde{U(1)}$, hence $R_{\xi}\circ|\cdot|\circ a\neq\tilde{R}_{\xi}$. Nevertheless, they coincide at the level of absolute values i.e.,
$(\cdot)^{\xi}\circ|\cdot|\circ a=|\cdot|^{r}\circ\tilde{R}_{\xi}=:|R_{\xi}|.$
The analogy with N-spectrum is partially explained by the following easy lemma.
Lemma 2.5.
Under Notation 1 and Definition 2.1 (ii), take any algebraic $\mathbb{C}$-scheme $X$. Then, any $\mathbb{C}$-morphism (as schemes) $\varphi\colon\Delta_{\xi}^{*,{\rm conv}}(\mathbb{C})\to X$ induces a holomorphic map $\varphi_{\epsilon}^{\rm an}\colon\widetilde{\mathbb{C}}_{\epsilon}^{*}\to X(\mathbb{C})$ for $0<\epsilon\ll 1$.
We leave the proof to readers as it is straightforward. Now, we generalize the notion of test configuration as follows. Henceforth, we stick to non-log setup i.e., of (often polarized) varieties without boundary divisors, as we leave the natural extensions to log pairs, which is very often almost verbatim now, to readers if necessary.
2.2. Generalized test configurations and examples
Henceforth, we always fix a base field $\mathbbm{k}$ throughout the paper (which is sometimes assumed to be $\mathbb{C}$ if mentioned). The following concepts implicitly or explicitly (compare [HL20, §2.2], [BJ21, Appendix A3], [Ino22, §2.2]) appeared in the literature. We introduce the terms just for slight extensions and convenience for our later usage, without any originality.
Definition 2.6.
  1. ($\mathrm{i}$)
    A generalized (unpolarized affine) test configuration (resp., generalized (polarized, projective) test configuration) is the following set of data:
    1. (a)
      free abelian groups $N$ of rank $r$.
    2. (b)
      algebraic tori $T$ defined as $T=N\otimes_{\mathbb{Z}}\mathbb{G}_{m}$.
    3. (c)
      a $r$-dimensional rational polyhedral cone $\tau(\supset\{0\})$ of $N_{\mathbb{R}}:=N\otimes\mathbb{R}$. and its corresponding affine toric variety $U_{\tau}$ over $\mathbbm{k}$. We denote the 00-dimensional strata of $U_{\tau}$ as $p_{\tau}$ or $0=p_{\tau}$ for simplicity.
    4. (d)
      a variety $\mathcal{Y}$ (resp., a polarized variety $(\mathcal{Y},\mathcal{L})$) and a $T$-action on it.
    5. (e)
      faithfully flat $T$-equivariant affine morphism $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$ (resp., $T$-equivariant projective flat morphism $p\colon(\mathcal{Y},\mathcal{L})\twoheadrightarrow U_{\tau}.$)
    6. (f)
      In the context when general fibers are affine cones (in the sense of [Od24a]), with respect to an additional algebraic $\mathbbm{k}$-torus $T^{\prime}$, we require more on the generalized (affine) test configuration as follows. $\mathcal{Y}$ admits $p$-fiberwise $T^{\prime}$-action so that it can be written as $\mathop{\mathrm{Spec}}\nolimits\mathcal{R}$ with a $\Gamma(U_{\tau})$-algebra $\mathcal{R}$ which we decompose as $\mathcal{R}=\oplus_{\vec{m}\in M}\mathcal{R}_{\vec{m}}$ via the $T^{\prime}$-action to the $T^{\prime}$-eigensubmodules over $\Gamma(U_{\tau})$, then each $R_{\vec{m}}$ is locally free over $\Gamma(U_{\tau})$.
  2. ($\mathrm{ii}$)
    For an affine integral reduced scheme $Y$ of finite type over a field $\mathbbm{k}$ (resp., polarized variety $(Y,L)$ over $\mathbbm{k}$), an (affine) algebraic test deformation (resp., (polarized) algebraic test deformation of $(T\curvearrowright Y,\xi)$), where $T$ is an algebraic $\mathbbm{k}$-torus $N\otimes\mathbb{G}_{m}$ i.e., which has the co-character lattice $N$, and $\xi$ is an element of $N\otimes\mathbb{R}$, is a generalized (unpolarized affine) test configuration $p\colon\mathcal{Y}\to U_{\tau}$ with $p^{-1}(0)=Y$.
    Note that if $\mathbbm{k}$ contains $\mathbb{R}$ and $T$ is $\mathbbm{k}$-split, $\xi$ can be identified with its holomorphic444we take this choice rather than real vector field, as in usual choice of Reeb vector field, as we want to work also over $\mathbbm{k}\neq\mathbb{C}$ vector field $\xi$ over the smooth locus $Y^{\rm sm}$ of $Y$. Indeed, if we consider the induced action of $T$ on $p^{-1}(0)$ and the natural map $a\colon N_{\mathbb{R}}\to H^{0}(p^{-1}(0),T_{p^{-1}(0)^{\rm sm}})$ induced by it (here, $T_{p^{-1}(0)^{\rm sm}}$ denotes the Zariski tangent sheaf of the smooth locus of $p^{-1}(0)$), the image ${\rm Im}(a)$ contains $a(\xi)$ and $a$ is often injective, so that we identify $\xi$ with a holomorphic vector field.
  3. ($\mathrm{iii}$)
    An affine algebraic test deformation of $(Y,\xi)$ over $U_{\tau}$ is algebraic positive (resp., negative) weight deformation (or $Y$ is algebraic scale up (resp., scale down) deformation) over $U_{\overline{\tau}}$ if
    1. (a)
      $T\curvearrowright Y$ is a good action in the sense of [LS13] (e.g., Fano cone as in [CS18, CS19]) and
    2. (b)
      $-\tau\ni\xi$ (resp., $\tau\ni\xi$).  [^5] Our sign convention comes from [Pin74, CH22] and matches with differential geometric intuition, while being opposite with some other algebraic literature.
    We also propose alternative names of algebraic positive (resp., negative) weight deformation: scale up (resp., scale down) degeneration, motivated by the differential geometric context.
Example 2.7.
Underlying family of affine test configuration (cf., [CS18]) is a generalized (unpolarized) test configuration with $r=1,\tau=\mathbb{R}_{\geq 0}$. Similarly, (polarized projective) test configuration ([Don02]) is a generalized (polarized projective) test configuration with $r=1,\tau=\mathbb{R}_{\geq 0}$.
Example 2.8.
If we consider an embedded affine variety $Y^{\prime}\subset\mathbb{A}_{\mathbbm{k}}^{l}$ and the natural action of $T=\mathbb{G}_{m}^{l}$ on the ambient space $\mathbb{A}_{\mathbbm{k}}^{l}$ (with all weights $1$). Then we consider $Y^{\prime}\times\{(1,\cdots,1))\}\subset\mathbb{A}_{\mathbbm{k}}^{l}\times T$ and its orbit $T\cdot(Y^{\prime}\times\{(1,\cdots,1))\})$. If the orbit closure $\overline{T\cdot(Y^{\prime}\times\{(1,\cdots,1))\})}$ inside $\mathbb{A}_{\mathbbm{k}}^{l}\times U_{\tau}$ is faithfully flat over $U_{\tau}$ for some (not necessarily strictly) convex rational polyhedral cone $\tau$ of $N_{\mathbb{R}}$, then
$\overline{T\cdot(Y^{\prime}\times\{(1,\cdots,1))\})}\twoheadrightarrow U_{\tau}$
gives a generalized (unpolarized) test configuration. This generalizes the classical construction of test configuration à la Mumford ([Mum65]) and Donaldson ([Don02]) for the $r=1$ case. However, the required (faithful) flatness becomes nontrivial in general for $r>1$.
Example 2.9 (The degenerations of klt singularities to its K-semistable Fano cone).
We discuss an algebraic version of the local version of the Donaldson-Sun theory [DS17]. Suppose the base field $\mathbbm{k}$ is algebraically closed of characteristic 00. For any affine klt variety $X$ with its closed point $x$, there is a canonical quasi-monomial valuation $v_{X}$ with the center $x$ ([Li18, Blu18]) and $W=\mathop{\mathrm{Spec}}\nolimits\mathop{\mathrm{gr}}\nolimits_{v_{X}}\mathcal{O}_{X,x}$ is a K-semistable Fano cone, in particular $\mathop{\mathrm{gr}}\nolimits_{v_{X}}\mathcal{O}_{X,x}$ is finitely generated ([LX18, XZ22]). Here, recall that $\mathop{\mathrm{gr}}\nolimits_{v_{X}}\mathcal{O}_{X,x}:=\oplus_{a\in\mathbb{R}_{\geq 0}}\{f\in\mathcal{O}_{X,x}\mid v_{X}(f)\geq a\}/\{f\in\mathcal{O}_{X,x}\mid v_{X}(f)>a\}$.
Choose its homogenous generators $\overline{z}_{1},\cdots,\overline{z}_{l}$, and we write their weights as $w_{1},\cdots,w_{l}\in\mathbb{R}_{>0}$ in the holomorphic spectrum (cf., [DS17]) and denote $\xi:=(w_{1},\cdots,w_{l})$. Naturally, we have the corresponding embedding $W\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}$. We then take lifts of $\overline{z}_{1},\cdots,\overline{z}_{l}$ to the element of $\Gamma(\mathcal{O}_{X})$, say $z_{1},\cdots,z_{l}\in\Gamma(\mathcal{O}_{X})$. By taking additional $z_{i}$s (and $\overline{z}_{i}$s) if necessary, we can assume that the corresponding map $X\hookrightarrow\mathbb{A}_{\mathbbm{k}}^{l}$, which maps $x$ to the origin $\vec{0}$, is an embedding.
Recall that by a slight perturbation of $v_{X}$ to a divisorial valuation $v^{\prime}$, we can easily show there is a natural isomorphism of the graded ring as $gr_{v_{X}}(\mathcal{O}_{X,x})=gr_{v^{\prime}}(\mathcal{O}_{X,x})$ (cf., e.g., [LX18, 2.10]) and we take associated test configuration as the spectrum of graded Rees algebra ([Mat89, Chapter 5, section 4]), which degenerates $x\in X$ to $W$ with the vertex. Note that this test configuration depends on the choice of the perturbation $v^{\prime}$. We integrate these degenerations for various approximation $v^{\prime}$ as follows.
For that, we first observe the following. We leave the proof to the readers as it is easy.
Lemma 2.10.
For any $\xi=(w_{1},\cdots,w_{l})\in\mathbb{R}^{l}$, we take a lattice $N$ which is the minimum sublattice of $\mathbb{Z}^{l}$ such that $N\otimes\mathbb{R}(\subset\mathbb{R}^{l})$ which contains $\mathbb{R}\xi$. Then the natural pairing map
$\langle\xi,-\rangle\colon M\to\mathbb{R},$
which sends $\vec{m}=(m_{1},\cdots,m_{l})\in M:={\rm Hom}(N,\mathbb{Z})$ to $\langle\vec{m},\xi\rangle=\sum_{i}m_{i}w_{i}$, is a group isomorphism onto its image $\Gamma:=\sum_{i}\mathbb{Z}w_{i}\subset\mathbb{R}$.
We apply the above for our weight vector $\xi=(w_{1},\cdots,w_{l})\in\mathbb{R}^{l}$, and also take a rational polyhedral cone $\tau$ of $N$ which contains $\xi$, to later discuss its perturbations inside $\tau$.
We define the algebraic subtorus $T\subset\mathbb{G}_{m}^{r}$ as $T:=\mathbb{G}_{m}\otimes N$. We take generators of the defining ideal of $X\subset\mathbb{A}_{\mathbbm{k}}^{l}$, say $f_{1},\cdots f_{a}$, which are even universal Gröbner basis of the defining ideal of $X$ ([BCRV22, §5.1]).
We then take small enough $\tau$ which includes $\xi$ and applies Example 2.8 which results in a flat $T$-equivariant family $\pi_{\tau}\colon\mathcal{X}_{\tau}\twoheadrightarrow U_{\tau}$ whose fiber is $X$ over $(1,\cdots,1)$ and the fiber over the $T$-invariant point $p_{\tau}$ is $W$ as the proof [LX18, 2.10] shows as $\tau$ is small enough.
By using the fact that $f_{i}$s form a universal Gröbner basis, it easily follows that for any rational ray $\mathbb{R}_{>0}\xi^{\prime}$ in $\tau$, the inducing valuation on $X$ is represented by $\mathbb{R}_{>0}\xi^{\prime}$, which we denote as $v_{\xi^{\prime}}$, ts extended Rees construction ([Mat89, Chapter 5, section 4]) is nothing but the restriction of $\pi_{\tau}$ to the orbit closure $\overline{(\mathbb{G}_{m}\otimes(N\cap\mathbb{R}\xi^{\prime}))\cdot[X]}$ along $\xi^{\prime}$-direction. Further, from the Zariski openness of the irreducibility condition, by replacing $\tau$ by a smaller enough one if necessary, we can and do assume that all the fibers of $\pi_{\tau}$ are irreducible. (Even more is true, as [LX18, 2.10] again shows that the fibers over the toric boundary are all isomorphic.) Thus its inducing valuation on $X$ is represented by $\mathbb{R}_{>0}\xi^{\prime}$. Further, from the construction, the above generalized test configuration $\pi_{\tau}$ is a positive weight deformation (scale up degeneration) in the sense of Definition 2.6.
[…]
We end Example 2.9 by the following brief summary of our discussion.
Theorem 2.11.
For any affine klt variety $X$ with its closed point $x$, over an algebraically closed field $\mathbbm{k}$ of characteristic 00, consider the volume-minimizing valuation $v_{X}$ of $\mathbb{Q}$-rank $r$ in the sense of [Li18], and $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v_{X}}(\mathcal{O}_{X,x}))$.
There is an affine toric variety $U_{\tau}$ and an algebraic positive weight deformation over it
$\pi_{\tau}\colon\mathcal{X}_{\tau}\twoheadrightarrow U_{\tau},$
which restricts to $W\times\partial U_{\tau}\twoheadrightarrow\partial U_{\tau}$, over the toric boundary $\partial U_{\tau}:=U_{\tau}\setminus T$. Its general fibers are $X$.
See closely related studies in [Tei03, §2] and [LX18, the proof of 2.10] which follows loc.cit.
[…]
Remark 2.12.
As the general context of [Mum65], consider a reductive algebraic group $G$ and its action on a polarized variety $(M,L)$ over an algebraically closed field $\mathbbm{k}$. Take a semistable closed point $x$ of $X$.
Then, the orbit closure $\overline{G\cdot x}\subset M$ contains a unique $G$-orbit which is closed inside the semistable locus. We denote it by $G\cdot y$ with $y\in X$. By [IM65] (cf., also [Mum65, Chapter 2, §1]), there is a one parameter subgroup $\lambda\colon\mathbb{G}_{m}\to G$ with $\lim_{t\to 0}\lambda(t)\cdot x=y$ though it is far from unique. If we take a maximal torus of $G$ which contains the image of $\lambda$, then clearly we have $\overline{T\cdot x}\ni y$, with a little less isomorphism classes. If $M$ is a moduli of certain varieties, this will give us a generalized test configuration but we do not know its canonicity in general (since $T$ is unique only up to the conjugation).
Remark 2.13.
Let us consider isotrivial degenerations of K-semistable Fano cone $W$ to K-polystable Fano cone $C$, which first appeared in [DS17] where the $C$ is the metric tangent cone of the Gromov-Hausdorff limits of Kähler-Einstein manifolds, and later [LWX21] gave an algebraic proof to the fact that such isotrivial degeneration exists for any $W$ as an affine test configuration.
Recall that $C$ is given an action of algebraic split torus $T$ from the definition, and by the reductivity of $G:={\rm Aut}_{T}(C)$ ([DS17, Appendix], cf., also [Od24a]), which means the group of automorphisms of $C$ which commutes with the $T$-action, and the action of its centralizer to the multi-Hilbert scheme [HM04] $H$ as in [DS17]. From these facts, it is easy to see similar remark apply as above Remark 2.12.
Example 2.14 ($\mathbb{R}$-test configurations, Kähler-Ricci flow).
For a general polarized variety $(X,L)$ over $\mathbbm{k}$, there is a notion of $\mathbb{R}$-test configuration (cf., [HL20, BLXZ23] as originated in [CSW18, §3]) which we recall now.
Definition 2.15.
If we set $R_{m}:=H^{0}(X,L^{\otimes m})$ for $m\in\mathbb{Z}_{>0}$, and $R_{(X,L)}:=\oplus_{m\in\mathbb{Z}_{\geq 0}}R_{m}$, the filtration $\mathcal{F}^{\lambda}R_{m}$ of $R_{m}$ for $\lambda\in\mathbb{R}$ which satisfies the following is called filtration of $(X,L)$ (cf., [Sze15]).
  1. ($\mathrm{i}$)
    $\mathcal{F}^{\lambda}R_{m}$ is a $\mathbbm{k}$-vector subspace of $R_{m}$ for each $\lambda\in\mathbb{R},m\in\mathbb{Z}$.
  2. ($\mathrm{ii}$)
    (decreasingness) $\mathcal{F}^{\lambda^{\prime}}R_{m}\supset\mathcal{F}^{\lambda}R_{m}$ for any $\lambda^{\prime}\leq\lambda$.
  3. ($\mathrm{iii}$)
    (linearly boundedness) there is $e_{-}$ and $e_{+}$ such that $\mathcal{F}^{e_{-}m}R_{m}=R_{m}$ and $\mathcal{F}^{e_{+}m}R_{m}=0$ for any $m$.
  4. ($\mathrm{iv}$)
    (multiplicativity) $\mathcal{F}^{\lambda_{1}}R_{m_{1}}\cdot\mathcal{F}^{\lambda_{2}}R_{m_{2}}\subset\mathcal{F}^{\lambda_{1}+\lambda_{2}}R_{m_{1}+m_{2}}$ for any $\lambda_{1},\lambda_{2}\in\mathbb{R},m_{1},m_{2}\in\mathbb{Z}$.
  5. ($\mathrm{v}$)
    (as a convention: left-continuous) $\cap_{\lambda^{\prime}<\lambda}\mathcal{F}^{\lambda^{\prime}}R_{m}=\mathcal{F}^{\lambda}R_{m}$.
For a filtration $\mathcal{F}^{\lambda}R_{m}$, we can define the graded ring
${\rm gr}_{\mathcal{F}}(R_{(X,L)})$ $:=\oplus_{m\in\mathbb{Z}_{\geq 0}}{\rm gr}_{\mathcal{F}}R_{m}$
$=\oplus_{m\in\mathbb{Z}_{\geq 0},\lambda\in\mathbb{R}}(\mathcal{F}^{\lambda}R_{m}/\cup_{\lambda^{\prime}>\lambda}\mathcal{F}^{\lambda^{\prime}}R_{m}).$
A filtration is called a $\mathbb{R}$-test configuration if ${\rm gr}_{\mathcal{F}}(R_{(X,L)})$ is finitely generated algebra over $\mathbbm{k}$.
Now we briefly recall the following geometric realization of $\mathbb{R}$-test configuration as a generalized test configuration, as essentially proved in [Tei03, HL20, BlJ20, Ino22] under the name of “geometric $\mathbb{R}$-test configuration II” ([HL20, p.10]), “rank $r$ test configuration” ([BlJ20, Appendix A.3]), and “polyhedral configuration” ([Ino22, §2.2.2, Prop 2.20]) respectively. See also loc.cit for more information on this correspondence.
Proposition 2.16 ([HL20, BlJ20, Ino22]).
For a $\mathbb{R}$-test configuration $\mathcal{F}$ of $(X,L)$, with its central fiber $(Y:={\rm Proj}({\rm gr}_{\mathcal{F}}(R(X,L))),\mathcal{O}(1))$ and its natural holomorphic vector field $\xi$, we can associate a canonical (polarized) algebraic test deformation of $(Y,\mathcal{O}(1),\xi)$ over a certain affine toric variety $U_{\tau}$.
Proof.
We proceed similarly as the proof of Theorem 2.11 in Example 2.9. Take homogeneous generators of ${\rm gr}_{\mathcal{F}}(R_{(X,L)})$ and their $\mathbb{Z}$-homogeneous lifts to $R(X,L)$ as $f_{1},\cdots,f_{l}$ and denote their gradings as $(m_{i},w_{i})(i=1,\cdots,l)$ with $m_{i}\in\mathbb{Z}_{\geq 0},w_{i}\in\mathbb{R}$. We set $\xi:=(w_{1},\cdots,w_{l})$ and take $\tau\ni\xi$ similarly. Then we apply Example 2.9 to the cone $Y^{\prime}:={\rm Spec}R(X,L)$ with the natural $\mathbb{G}_{m}$-action. By dividing the resulting family by $\mathbb{G}_{m}$, we obtain the desired family. We omit the details as the proof is almost verbatim to Example 2.9. ∎
[…]
Recall that [CW20, Theorem 1.6] and then [Bam18, Corollary 1.3], [WZ21, Theorem 1.1] proved the Hamilton-Tian conjecture which says that the Gromov-Hausdorff limit $W$ of the normalized Kähler-Ricci flow on (possibly K-unstable) Fano manifold $X$ is $\mathbb{Q}$-Fano variety with weak Kähler-Ricci soliton, defined as
$\frac{\partial}{\partial t}\omega_{t}=\omega_{t}-{\rm Ric}(\omega_{t}),$
with $[\omega_{t}]\in 2\pi c_{1}(X).$ Then [CSW18] proved even more i.e., it is obtained as the central fiber of $\mathbb{R}$-test configuration of $(X,\mathcal{O}(-K_{X}))$, which minimizes the H-invariant as proved in [DS20, HL20] (cf., also [BLXZ23]).
Corollary 2.17 (cf., [CSW18, HL20, BLXZ23]).
For any K-unstable $\mathbb{Q}$-Fano variety $X$, consider the corresponding a (klt) $\mathbb{Q}$-Fano variety $W$ with action of an algraic torus $T$ and $\xi\in({\rm Hom}(\mathbb{G}_{m},T)\otimes\mathbb{R})\setminus({\rm Hom}(\mathbb{G}_{m},T)\otimes\mathbb{Q})$, which satisfies the (modified) K-semistability in the sense of [BN14, §3.5], [HL20, §2]. This is the $1$-st process of the degeneration to the limit of the normalized Kähler-Ricci flow on $X$ as [CSW18, HL20], also characterized as the linear bounded filtration whose $H^{\rm NA}$-invariant is minimized ([HL20, BLXZ23]). Thus, there is a canonical algebraic test deformation of $W$ to $X$.
Example 2.18 (Negative weight deformations).
Conlon-Hein [CH22] introduces a notion of negative weight deformation of Calabi-Yau cones in terms of sequence of affine test configurations, for their classification of asymptotically conical Calabi-Yau manifolds. We alternatively introduce the notion of algebraic negative deformation in Definition 2.6 without using sequence (to answer Open Questions 1.4(2) of [CH22]). For our later purpose, we also clarify the relation as follows. Here, we assume $\mathbbm{k}=\mathbb{C}$ as [CH22]. We take a affine $\mathbb{C}$-variety $Y$ and an algebraic split torus $T=N\otimes\mathbb{G}_{m}$ acting on $Y$, with $\xi\in N_{\mathbb{R}}$.
Lemma 2.19 (Rational approximation - relation with [CH22]).
  1. ($\mathrm{i}$)
    Suppose $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$ is an algebraic negative weight deformation (Definition 2.6) of $(Y,\xi)$. In particular, $\xi\in\tau$. Then, its general fiber is a negative $\xi$-weight deformation in the sense of [CH22, Definition 1.7].
  2. ($\mathrm{ii}$)
    ([Od24c]) Conversely, if $Y$ has only isolated singularities and $T$ acts as a good action in the sense of [LS13], any negative $\xi$-weight deformation of normal affine variety with $Y$ and $\xi$ arise in this way.
Proof.
For the item ($\mathrm{i}$), we can assume that $\tau=\mathbb{Z}_{\geq 0}^{r}$, $N=\mathbb{Z}^{r}$. Then we approximate $\xi$ by $\xi_{i}=(\frac{r_{i,1}}{m_{1}},\frac{r_{i,2}}{m_{1}},\cdots,\frac{r_{i,r}}{m_{1}})\in\tau\cap N_{\mathbb{Q}}$, with nonnegative integers $r_{i,j}(j=1,\cdots,r)$ and $m_{i}$s so that a convergence $\xi_{i}\to\xi$ holds for $i\to+\infty$. Then take a morphism $\varphi_{\vec{r}}\colon\mathbb{A}^{1}\to U_{\tau}$ defined by $t\mapsto(t^{r_{i,1}},t^{r_{i,2}},\cdots,t^{r_{i,r}})$ and consider the base change $\mathcal{Y}\times_{p,U_{\tau},\varphi_{\vec{r}}}\mathbb{A}^{1}$, which we denote as $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}_{1}$. It can be naturally seen as an affine test configuration but we modify the $\mathbb{G}_{m}$-action $\sigma^{\prime}_{i}$ on it by compositing with the $m_{i}$-th power map $\mathbb{G}_{m}\to\mathbb{G}_{m}$. In particular, $p^{\prime}_{i}(\sigma^{\prime}_{i}(z,x))=z^{m_{i}}p^{\prime}_{i}(x)$ for $z\in\mathbb{G}_{m}(\mathbbm{k}),x\in W^{\prime}_{i}$. Then, we take the base change of $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}^{1}$ with $\sigma^{\prime}_{i}$ by the $m_{i}$-th power map $q_{i}\colon\mathbb{A}^{1}\to\mathbb{A}^{1}$ whose target is regarded as the base of $p^{\prime}_{i}$. It is straightforward confirmation that the sequence, with respect to $i=1,2,\cdots$, of the obtained fiber product $p_{i}\colon W_{i}(:=W^{\prime}_{i}\times_{p^{\prime}_{i},\mathbb{A}^{1},q_{i}}\mathbb{A}^{1})\to\mathbb{A}^{1}$ with the $\mathbb{G}_{m}$-action induced by $\sigma^{\prime}_{i}$ satisfies the conditions of the [CH22, Definition 1.7], with $c_{i}$ of loc.cit as $m_{i}$, $\mu_{i}$ of loc.cit as $1$, hence the latter being auxiliary parameters. Suppose $Z$ is the maximum Artin scheme supported at the $T$-invariant point $p_{\tau}\in U_{\tau}$, such that $\mathcal{Y}|_{Z}$ is trivial family. Such $Z$ exists due to the Zorn’s lemma combined with the formal smoothness ([Gro64, Part I, Part IV]) of ${\rm Aut}_{T}(Y)$ which is an algebraic group as the stabilizer of $[Y]$ in the multi-Hilbert scheme. By using this $Z$, $k_{i}$ of loc.cit can be written as $m_{i}{\rm length}\mathcal{O}_{\varphi_{\vec{r}}^{*}Z}$, which explains the nature of the $\xi$-weight in loc.cit. ∎
The item ($\mathrm{ii}$) is proved in [Od24c] in which some more related materials are also explored.
3. Higher rank semistable reduction theorem
In this section, we generalize the theorem of [AHLH23, §6] to the families over higher dimensional toroidal base or Novikov type rings.
3.1. Toroidal stack-theoretic setup
Here are some preparations on stack-theoretic framework, following [HL14, Hei18, AHLH23].
Definition 3.1.
Under the notation 1 in the previous subsection, we set the following Artin stacks.
  1. ($\mathrm{i}$)
    Consider the quotient stack of the affine toric variety $U_{\tau}$:
    $\Theta_{\tau}:=[U_{\tau}/T],$
    generalizing the case $r=1,\tau=\mathbb{R}_{\geq 0}$ treated in [HL14, AHLH23], which is denoted simply as $\Theta$. We denote the $T$-invariant point of $U_{\tau}$ as $p_{\tau}$. In [HL14], (the isomorphic object) $\Theta_{\tau}$ for the case of regular simplicial $\tau$ is also essentially introduced under the disguise of $\Theta^{r}$, the $r$-times self-product of $\Theta$. For our purpose, since we will encounter a setup with various $\tau$ for a fixed $N$ later, we prefer the above setup and notation.
  2. ($\mathrm{ii}$)
    We fix a valuation ring $(R,v)$ of height $1$ with $\Gamma_{\geq 0}:=v(R\setminus\{0\})\subset\mathbb{R}$. To set up a notion of elementary modification of family along $R$ (Definition 3.4 later), we prepare the following.
    Further, we set $N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$ and take a $r$-dimensional polyhedral cone $\tau^{\prime}$ inside $N_{\mathbb{R}}\oplus\mathbb{R}$, whose generators can be taken inside $N\times\Gamma_{\geq 0}$ and assume that the restriction of the first projection $N_{\mathbb{R}}\times\mathbb{R}\to N_{\mathbb{R}}$ to $\tau^{\prime}$ is a PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$. In other words, $\tau^{\prime}$ is the graph of a homomorphism $\tau\to\mathbb{R}_{\geq 0}$.
    We further assume the normal vector of $\tau$ can be taken as $(\vec{m},1)\in M\oplus(\Gamma_{\geq 0})^{*}$, i.e., $\langle\tau,(\vec{m},1)\rangle=0$. We denote the character on $T$ which corresponds to $\vec{m}$ as $\chi_{\vec{m}}$. Then we define
    (4) $\widetilde{\rm ST}_{R}(\tau^{\prime})$ $:=\mathop{\mathrm{Spec}}\nolimits R[\tau^{\vee}\cap M]\biggl{(}\frac{\pi}{\chi_{\vec{m}}}\biggr{)},$
    (5) ${\overline{\rm ST}}_{R}(\tau^{\prime})$ $:=[\widetilde{\rm ST}_{R}(\tau^{\prime})/T],$
again generalizing the case $r=1,\tau=\mathbb{R}_{\geq 0}$ introduced in [Hei18, §2.B] and [AHLH23, §3.5]. Note that, if $R=\mathbbm{k}[[t]]$, the adjunction of $\frac{\pi}{\chi_{\vec{m}}}$ in the right hand side of (4) is the base change of a localization corresponding to a toric blow up of $R\times U_{\tau}$, with respect to
(6) $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1).$
In general, we have a natural (birational) map
$b\colon{\overline{\rm ST}}_{R}(\tau^{\prime})\to{\rm Spec}(R)$
and a canonical open embedding ${\rm Spec}(K)\hookrightarrow{\overline{\rm ST}}_{R}(\tau^{\prime}),$ where $K$ denotes the fraction field of the valuation ring $R$. We denote the one dimensional strata of $\widetilde{\rm ST}_{R}(\tau^{\prime})\simeq\mathbb{G}_{m}(\mathbbm{k})$ corresponding to $\tau^{\prime}$ (or its image in ${\overline{\rm ST}}_{R}(\tau^{\prime})$) as $V(\tau^{\prime})$.
Definition 3.2 (Complex analytic analogue).
If $\mathbbm{k}=\mathbb{C}$, then we can literally make sense of the toroidality of the blow up $b^{\rm an}\colon U_{\hat{\tau}^{\prime}}\to\mathbb{C}\times U_{\tau}^{\rm an}$. We denote the complex analytic space $(b^{\rm an})^{-1}(\Delta\times U_{\tau}^{\rm an})$ as $\widetilde{\rm ST}_{\Delta}^{\rm an}(\tau^{\prime})$ its analytic quotient stack ${\overline{\rm ST}}_{\Delta}^{\rm an}(\tau^{\prime}):=[\widetilde{\rm ST}_{\Delta}^{\rm an}(\tau^{\prime})/T(\mathbb{C})^{\rm an}]$.
Now we extend the notion of $\Theta$-strata ([HL14]) and elementary modifications ([AHLH23]) accordingly, which morally corresponding to the “optimal destabilization”. The point of our generalization is that they are not necessarily one parameter algebraic families but rather higher dimensional families, which contains “irrational” directions.
Below, we fix an algebraic stack $\mathcal{M}$ over $\mathbbm{k}$.
Definition 3.3 (Higher $\Theta$-strata cf., [HL14, §2.1] for $r=1$ case).
Consider the mapping stack ${\rm Map}(\Theta_{\tau},\mathcal{M})$ of $\Theta_{\tau}$ to $\mathcal{M}$, as an algebraic stack, after [AHR20, 5.10, 5.11], [AHR19, 6.23], [HLP23, 5.1]. Extending [HL14, Definition 2.1.1], a higher $\Theta$-stratum of type $\tau$ and rank $r$ in $\mathcal{M}$ consists of a union of connected components $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$ such that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed embedding. Note that, if $\tau$ is a regular simplicial one and we fix an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for a rational polyhedral cone $\tau$, we can regard it as an open closed substack of ${\rm Filt}^{r}(\mathcal{M}):={\rm Map}(\Theta^{r},\mathcal{M})$.
Furthermore, if $\mathcal{M}$ is quotient stack of the form $[M/G]$ where $M$ is an algebraic $\mathbbm{k}$-scheme $M$ and $G$ is an algebraic $\mathbbm{k}$-group, we often write ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=[Z^{+}/G]$ for some $G$-invariant closed subscheme $Z^{+}\subset M$.
Now, we are ready to introduce the generalization of elementary modification of [AHLH23, §3].
Definition 3.4 (Elementary ($\tau$-)modification cf., [AHLH23, §3.5, Definition 3.36] for $r=1$ case).
An extended elementary ($\tau$-)modification of $f\colon{\rm Spec}(R)\to\mathcal{M}$ is the data of a morphism $\overline{f}\colon\overline{\rm ST}_{R}(\tau)\to\mathcal{M}$ along with an isomorphism $f\simeq\overline{f}|_{\rm Spec(R)}$. Here, to make sense of the restriction symbol, $\rm Spec(R)$ is regarded as a localization of $\overline{{\rm ST}}_{R}(\tau^{\prime})$ since we have $\mathop{\mathrm{Spec}}\nolimits(R)=[\mathop{\mathrm{Spec}}\nolimits R[M]/T]$.
The following lemma lies the ring-theoretic aspect of Definition 2.3 and it appears in Theorem 3.8 later.
Lemma 3.5 (Extension of valuation ring).
Suppose $(R,v=v_{R})$ is a valuation local ring, such that the associated graded ring $gr_{v}(R):=\oplus_{w\in{\rm Im}(v)}\{a\in R\mid v(a)\geq w\}/\{a\in R\mid v(a)>w\}$ is isomorphic to a semigroup ring over a field $\mathbbm{k}$ and a submonoid of $\mathbb{R}^{r}$ with some $r\in\mathbb{Z}_{>0}$ as a graded ring (which holds e.g. if $R$ is of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ for a field $\mathbbm{k}$ in Definition 2.1) and $\xi,N,M_{\xi\geq 0}$ are as Notation 1. Now, we define a ring $R[[M_{\xi\geq 0}]]$ as
$\{\sum_{i=1,2,\cdots}a_{i}t^{m_{i}}\mid a_{i}\in R,m_{i}\in M_{\xi\geq 0},m_{1}<m_{2}<\cdots,m_{i}\to\infty(i\to\infty)\},$
as Definition 2.1 (i), and consider
$R_{\rm new}(\xi)^{\prime}:=R[[M_{\xi\geq 0}]]/(\pi-\chi^{m}),$
for $(0\neq)\pi\in R\setminus R^{*}$ and $m\in M_{\xi\geq 0}$ which we suppose to be primitive i.e., there is no positive integer $a>1$ with $m^{\prime}\in M_{\xi\geq 0}$ such that $am^{\prime}=m$.
Then, $R_{\rm new}(\xi)^{\prime}$ has a natural valuation $\overline{v}_{\rm new}$ induced from $v_{R}$, with $\overline{v}_{\rm new}(\overline{\pi})=v(\pi)$. We take its associated valuation ring $R_{\rm new}(\xi)$ in the fraction field of $R_{\rm new}(\xi)^{\prime}$. Then, the associated graded ring is again isomorphic to a semigroup ring over a field, for a submonoid of $\mathbb{R}^{r}$ with some $r\in\mathbb{Z}_{>0}$ as a graded ring.
Proof.
Set $b:=\frac{v(\pi)}{\langle\xi,m\rangle}>0$ and put $\tilde{v}(\sum a_{i}t^{m_{i}}):=\min_{i}\{v(a_{i})+b\langle m_{i},\xi\rangle\}$ for $a_{i}\in R,m_{i}\in M_{\xi\geq 0}$. This $\tilde{v}$ defines a valuation on $R[[M_{\xi\geq 0}]]$ and consider its corresponding filtration $\mathcal{F}_{\tilde{v}}$ as $\mathcal{F}_{\tilde{v}}(x)=\{f\in R[[M_{\xi\geq 0}]]\mid\tilde{v}(f)\geq x\}$ for each $x\in\mathbb{R}$. It induces a filtration $\overline{\mathcal{F}}$ on $R_{\rm new}(\xi)$ as a quotient of $R[[M_{\xi\geq 0}]]$, and the corresponding graded ring is clearly $({\rm gr}_{v}(R))[M_{\xi}/\mathbb{Z}_{\geq 0}m]$ which is still a semigroup over $\mathbbm{k}$ ring for a submonoid of $\mathbb{R}^{r}$ for a bigger $r$. Since it is an integral domain, $\overline{\mathcal{F}}$ also induces a valuation on $R_{\rm new}(\xi)$. ∎
Remark 3.6 (Novikov-type case).
Even if $R$ is of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ in Definition 2.1, hence a valuation ring, $R_{\rm new}(\xi)^{\prime}$ is not necessarily of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ for a different $\Gamma_{\geq 0}$ as in Definition 2.1, and it is not a valuation ring in general. For instance, such a pathology often happens if we continue the above process $R\rightsquigarrow R_{\rm new}(\xi)$ for the same irrational $\xi\in M_{\mathbb{R}}$ twice. Nevertheless, if $R$ is of such form $\mathbbm{k}[[\Gamma_{\geq 0}]]$, $R_{\rm new}(\xi)$ is again of such form $\mathbbm{k}[[\Gamma_{\geq 0}+bM_{\xi\geq 0}]]$ for $b\in\mathbb{R}_{>0}$ as in the proof of Lemma 3.5.
Remark 3.7 (Relation with $\overline{\Delta_{\xi}}(R)$).
In the setting of Definition 2.3, a formal completion of the inductive limit of $R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})$ is nothing but $R_{\rm new}(\xi^{\prime})$ for $\xi^{\prime}\in N^{\prime}\otimes\mathbb{R}$ with some $m\in M^{\prime}$ defined in the above Lemma 3.5. (The place of is not a typo.) Hence, $R_{\rm new}(\xi^{\prime}(=(\xi,c))$ can be understood as the completion of a stalk of $\overline{\Delta_{\xi}}(c,\pi,R)$ of Definition 2.3.
3.2. Higher $\Theta$-semistable reduction theorem
Below, we generalize the semistable reduction type theorem in [AHLH23, §6.1], which was for $r=1$ case. Morally speaking, it says that a one parameter family which strictly degenerates into a higher $\Theta$-strata, there is a canonical modification of the family to avoid a higher $\Theta$-strata i.e., with “less degeneration”. This is a version of semistable reduction type theorem, and later Theorem 3.18 is its iterated version along a certain kind of stratification (higher $\Theta$-stratification cf., Definition 3.17).
Theorem 3.8.
Let $\mathbbm{k}$ be a field, $(R,\mathfrak{m},v)$ be an equicharacteristic divisorial valuation ring which is $\mathbbm{k}$-algebra with $R/\mathfrak{m}\simeq\mathbbm{k}$ (divisorial valuation $\mathbbm{k}$-algebra), or more generally a valuation $\mathbbm{k}$-algebra which satisfies Lemma 3.5, and $\mathcal{M}$ be an algebraic stack over $\mathbbm{k}$ which is a quotient stack $[M/G]$ where $M$ is an algebraic scheme over $\mathbbm{k}$ and $G$ is an algebraic affine group acting on $M$.
On the other hand, as in Notation 1, we set $T=\mathbb{G}_{m}\otimes_{\mathbb{Z}}N$, where $N$ is a free abelian group of rank $r$, and fix a rational polyhedral cone $\tau\subset N_{\mathbb{R}}$ and $(0\neq)\xi\in\tau$ which is irrational if $r>1$.
Consider an arbitrary morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$. Suppose that $f(\mathfrak{m})$ sits in a higher $\Theta$-stratum $\mathcal{Z}^{+}$ in $\mathcal{M}$ while $f(\eta_{R})$ is not in $\mathcal{Z}^{+}$, where $\eta_{R}$ denotes the generic point of $R$. Now we fix some $\xi\in\tau\setminus N\otimes\mathbb{Q}$. Then, after
  • replacing $\tau$ by a smaller rational poyhedral cone of $N_{\mathbb{R}}$ which still contains $\xi$ (if necessary)
  • replacing $R$ by another divisorial valuation $\mathbbm{k}$-algebra with local $\mathbbm{k}$-morphism $R\to R^{\prime}$, so that the valuation of $R^{\prime}$ restricts to that of $R$.
  • taking a rational polyhedral cone $\tau^{\prime}$ of $N\oplus\mathbb{Z}$ which projects as PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$, as in Definition 3.1,
there exists the following canonical extension $\overline{f}$ of $f$:
  1. (i)
    $f$ extends to its unique (for fixed $\tau$) extended elementary $\tau$-modification (see Definition 3.4)
    (7) $\overline{f}\colon{\overline{\rm ST}}_{R}(\tau^{\prime})\to\mathcal{M}$
    which satisfy the following (note that it is an extension of $f$):
    The generic point $\eta_{V(\tau^{\prime})}$ of the one dimensional locus $V(\tau^{\prime})$ (recall from Definition 3.1) maps to outside $Z^{+}$ and there is a $T$-equivariant morphism from the N-spectrum $\iota\colon\Delta_{-\xi}\to\mathcal{M}$ which satisfies that
    • $\overline{f}(\eta_{V(\tau^{\prime})})\notin Z^{+}$ (avoiding the higher $\Theta$-strata),
    • $\iota(\eta_{-\xi})=\overline{f}(\eta_{V(\tau^{\prime})}),$ where $\eta_{-\xi}$ denotes the generic point of the N-spectrum $\Delta_{-\xi}$,
    • $\iota(c_{-\xi})=\overline{f}(\eta_{V(\hat{\tau}^{\prime})})\in Z^{+}$, where $c_{-\xi}$ denotes the closed point of the N-spectrum $\Delta_{-\xi}$.
  2. (ii)
    For any $\xi^{\prime\prime}\in(N\otimes\mathbb{Q})\cap\tau$, if we consider the one parameter subgroup $\lambda_{\xi^{\prime\prime}}\colon\mathbb{G}_{m}\to T$, then we have $\lim_{t\to 0}\lambda_{\xi^{\prime\prime}}(t)\cdot\iota(\eta_{-\xi})=\iota(c_{-\xi}).$
  3. (iii)
    (Canonicity of degenerations) Just in this item (iii), assume $\mathcal{M}$ is quotient stack by an algebraic split torus i.e., of the form $[M/T]$ where $M$ is an algebraic $\mathbbm{k}$-scheme and $T$ an algebraic split $\mathbbm{k}$-torus, and fix a lift of $f$ as $\tilde{f}\colon{\rm Spec}(R)\to M$. Then, the $\iota$ in (i), with admits some $\overline{f}$ which lifts to $\widetilde{\rm ST}_{R}(\tau^{\prime})$, is unique up to translation by $T(\mathbbm{k})\simeq(\mathbbm{k}^{*})^{r}$.
Remark 3.9.
If $\mathcal{M}$ is a moduli stack of varieties and $\overline{f}$ lifts to a morphism from $\widetilde{\rm ST}_{R}(\tau^{\prime})$ to a chart $M$ of $\mathcal{M}$, on which $T$ acts, its restrictions to each of the two strata corresponding to $\tau^{\prime}$ and $\tau$ gives a generalized test configurations in the sense of Definition 2.6.
By restricting the obtained extension $\bar{f}$ of $f$, we obtain the semistable reduction type statement, generalizing [AHLH23, §6].
Corollary 3.10 (Semistable reduction along higher $\Theta$-strata).
In the setup of Theorem 3.8, there exists a canonical modification $f|_{\xi^{\prime}}\colon\overline{\Delta_{\xi}}(R^{\prime})\to\mathcal{M}$ for $\xi^{\prime}=(\xi,c)\in\tau^{\prime}$ (cf., Definition 2.3) of $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ such that
  • it extends $f|_{\rm Spec(K)}\colon\mathop{\mathrm{Spec}}\nolimits(K)\to\mathcal{M}$,
  • the closed point $c_{\xi}$ maps outside of $\mathcal{Z}^{+}$,
  • $\lim_{t\to\infty}\lambda_{\xi^{\prime\prime}}(t)\cdot f|_{\xi^{\prime}}(c_{\xi})\in\mathcal{Z}^{+}$ for any $\xi^{\prime\prime}\in N\otimes\mathbb{Q}$ which approximates $\xi$ enough (as (ii)).
In particular, we obtain a canonical modification
$f_{{\rm new},\xi}\colon\mathop{\mathrm{Spec}}\nolimits R_{\rm new}(\xi)\to\mathcal{M}$
which maps the closed point to outside $\mathcal{Z}^{+}$ (recall Lemma 3.5).
Here, the Novikov type valuation spectrum/ring $\overline{\Delta_{\xi}}(R^{\prime})$ and its complete version $R_{\rm new}(\xi)$ are used to play the role of the new base, as it is natural (or canonical) from the construction in the actual proof below. Indeed, if we are satisfied with extension $R^{\prime}$ of $R$ as discrete valuation ring for $r>1$ case, there are so many choices (as various smaller sub $\mathbbm{k}$-algebras of $R_{\rm new}(\xi)$ or their completions etc).
proof of Corollary 3.10 (assuming Theorem 3.8).
If we consider the direction along $\tau^{\prime}$, note that $\overline{f}$ restricts to $[(\mathop{\mathrm{Spec}}\nolimits R[(\tau^{\prime})^{\vee}\cap M^{\prime}]/(\pi-p(\vec{0},1))/T]\to\mathcal{M}$ which maps the closed point to outside $\mathcal{Z}^{+}$, where $M^{\prime}$ is the dual lattice of the lattice $N^{\prime}$ which is the minimal sublattice of $N$ such that $N^{\prime}_{\mathbb{R}}$ contains $\xi^{\prime}$, and $(\tau^{\prime})^{\vee}$ denotes the dual cone. We denote the morphism as $\overline{f}|_{\tau^{\prime}}$, for simplicity.
Then, if we take the projective limit with respect to $\tau^{\prime}$, and take composite with the completion map (see Proposition 2.2) $f|_{\xi^{\prime}}$ of Corollary 3.10 is obtained. ∎
proof of Theorem 3.8.
Our proof of Theorem 3.8 is divided into steps. The core idea and arguments are in Step 1 with the key lemma 3.11, which may be one feature of this paper, and the rest is about how to reduce to it.
Step 1:
for quotient stack by an algbraic torus,
Step 2:
for general quotient stack (by an affine algberaic group)
Step 1.
First, we assume that $\mathcal{M}=[M/T]$ with a presentation $M$ and an algebraic split $\mathbbm{k}$-torus $T$ and prove the theorem in this case. Later, we essentially reduce to this case. From our assumption on the structure of higher $\Theta$-strata on $\mathcal{Z}^{+}$ and the Białynicki-Birula type theorem [HL14, 1.4.7], $\mathcal{Z}^{+}=[Z^{+}/T]$ for some closed subscheme $Z^{+}\subset M$ and there is a morphism $f\colon{\rm Spec}(R)\to M$ and $T$-equivariant morphism $Z^{+}\times U_{\tau}\to M$ which extends $f$, following Notation 1, for some rational polyhedral cone $\tau$. Since there is also the (compatible) $T$-action on $M$, they induce a $T$-equivariant rational map $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$. From the condition that $Z^{+}$ gives a union of connected components in the mapping stack, it follows that $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$ is indeed not defined in the neighborhood of $(p_{\tau},\mathfrak{m})$. We resolve its indeterminancy in an explicit “toroidal” manner as the follwing lemma.
This is our core for the proof of Theorem 3.8 and is of purely scheme-theoretic nature.
Lemma 3.11 (key lemma).
If necessary, by replacing $\tau$ by a smaller rational polyhedral cone which still contains $\xi$, there is a regular simplicial subdivision fan of finite rational polyhedral subcones of $\tau\times\mathbb{R}_{\geq 0}$, which we denote as $\Sigma^{\prime}$ such that the following holds:
[…]
  • There is some $c^{\prime}\in\mathbb{R}_{>0}$ such that $\mathbb{R}_{\geq 0}\xi^{\prime}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ with $\xi^{\prime}=(\xi,c^{\prime})$ lies inside a $r$-dimensional rational polyhedral cone $\tau^{\prime}$ of $\Sigma^{\prime}$ and the convex hull
    $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1)$
    is a cone inside $\Sigma^{\prime}$. Further, $\tilde{f}(V({\tau^{\prime}}),0)\notin Z^{+}$, where $V(\tau^{\prime})$ denotes the strata of $\mathcal{U}_{\Sigma^{\prime}}$ which corresponds to $\tau^{\prime}$ as in usual toric geometry (hence, it is an image of a $\mathbb{G}_{m}$-torsor).
proof of Lemma 3.11.
We consider $T$-equivariant local coordinates $z_{1},\cdots,z_{N}$ around the image of the $T$-invariant point $p_{\tau}\in U_{\sigma}$ and consider its pullback (as rational functions) to $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)$ and denote them as $\frac{f_{i}}{g_{i}}$ with $T$-semiinvariant holomorphic (algebraic) $f_{i}$s and $g_{i}$s. Consider the ideal $\mathcal{J}:=\langle\{{\rm lcm}\{g_{i}\},f_{1}\frac{{\rm lcm}\{g_{i}\}}{g_{1}},\cdots,f_{N}\frac{{\rm lcm}\{g_{i}\}}{g_{N}}\}\rangle$ which is $T$-invariant from the definition and its blow up resolves the indeterminancy of the rational map $\mathop{\mathrm{Spec}}\nolimits R\times U_{\tau}\to M$ from the definition. Here, ${\rm lcm}\{-\}$ stands for the least common multiple but any regular function on ${\rm Spec}(R)\times U_{\tau}$ which divides $g_{i}$s for all $i$ plays the role.
From the complete reducibility of $T$, $\mathcal{J}$ which can be decomposed as
(8) $\mathcal{J}=\sum I_{\vec{m}}z^{\vec{m}},$
where $\vec{m}$ runs over elements of $\tau^{\vee}\cap M$, $z^{\vec{m}}$ means $\prod_{i}z_{i}^{m_{i}}$, and $I_{\vec{m}}$ are ideals of $R$. With respect to the natural partial order on $\tau^{\vee}\cap M$ as $\vec{m}\leq\vec{m^{\prime}}:\Leftrightarrow\vec{m^{\prime}}-\vec{m}\in\tau^{\vee}\cap M$, we have $I_{\vec{m}}\subset I_{\vec{m}}$ if $\vec{m}\leq\vec{m^{\prime}}$. If we assume $R$ is a discrete valuation ring for simplicity, we take a uniformizer of $R$ as $\pi$ and we can write $I_{\vec{m}}$ as $(\pi^{l_{m}})$, with $l_{m}\in\mathbb{Z}_{\geq 0}$. In this case, we have $l_{\vec{m}}\geq l_{\vec{m^{\prime}}}\geq 0$ if $\vec{m}\leq\vec{m^{\prime}}$.
Note that a uniformizer $\pi$ of $R$ gives a formally étale morphism to an affine $\mathbbm{k}$-line 777this is where we use the equi-characteristic assumption of $R$, but we believe this is likely only for the simplicity (and geometric nature) of our description for now. A more appropriate algebraic translation of the arguments should directly extend to the mixed characteristics case. which we write $c_{\pi}\colon{\rm Spec}(R)\to\mathbb{A}^{1}_{\mathbbm{k}}$. From the form (8) of $\mathcal{J}$, there is a $T$-invariant ideal $J\subset\Gamma(\mathcal{O}_{U_{\tau}\times_{\mathbbm{k}}\mathbb{A}^{1}_{\mathbbm{k}}})$ whose pullback by $c_{\pi}$ coincides with $\mathcal{J}\subset\Gamma(\mathcal{O}_{U_{\tau}\times_{\mathbbm{k}}{\rm Spec}(R)})$. Because of the $T$-invariance of $J$, normalization of the blow up of $U_{\tau}\times\mathbb{A}^{1}$ along $J$ is a toric $\mathbbm{k}$-morphism with respect to the subdivision of $\tau\times\mathbb{R}_{\geq 0}$ to a rational polyhedral fan $\Sigma^{\prime}$ i.e.,
$b_{\Sigma^{\prime}}\colon T_{N}{\rm emb}(\Sigma^{\prime})\to\mathbb{A}^{1}_{\mathbbm{k}}\times_{\mathbbm{k}}U_{\tau},$
[…]
We can and do assume henceforth that $\Sigma^{\prime}$ is a regular simplicial fan, by taking finer subdivision if necessary. Suppose that the rays of $\Sigma^{\prime}|_{\mathbb{R}_{\geq 0}\xi+\mathbb{R}_{\geq 0}(\vec{0},1)}$ are spanned by $(\xi,0)$ and $\{(\xi,c_{i})\mid i=1,2,\cdots,l\}$ with $c_{1}>c_{2}>\cdots>c_{l}$ and denote the $r$-dimensional rational polyhedral cone which includes $(\xi,c_{i})$ as $\tau^{\prime}_{i}$ for $i\leq l$, and which includes $(\xi,0)$ as $\tau_{l+1}=\tau^{\prime}_{l+1}$.
By further subdividing $\Sigma^{\prime}$ and taking its subfan if necessary, we can and do assume that all $\tau^{\prime}_{i}$ projects down to the same cone $\tau$ in $N_{\mathbb{R}}$ which is regular simplicial, its support stil contains $\xi$, and that $\{\tau^{\prime}_{i}\}_{i}\cup\{\tau\times\{0\}\}$ are the only $r$-dimensional cones (walls) in $\Sigma^{\prime}$, and we also set $\tau_{l+1}=\tau\times\{0\}$. Now, we take a $(r+1)$-dimensional rational polyhedral cone of $\Sigma^{\prime}$ which includes the sub-interval
$(\xi,[c_{i}-\epsilon,c_{i}])\subset N_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$
for $0<\epsilon\ll 1$ as $\tilde{\tau_{i}}$ for each $i=1,\cdots,l$, and likewise, a $(r+1)$-dimensional rational polyhedral cone of $\Sigma^{\prime}$ which includes the sub-interval
$(\xi,[c_{1},c_{1}+\epsilon])\subset N_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$
as $\tilde{\tau}^{\prime}_{0}$. Also, just as a notation in this paper, we denote each $T$-orbit of $T_{N}{\rm emb}(\Sigma^{\prime})$ which corresponds to a cone $\sigma\in\Sigma^{\prime}$ as $\overline{V}(\sigma)$ and its pullback to $\mathcal{U}_{\Sigma^{\prime}}$ as $V(\sigma)$.
Note that $V(\tilde{\tau_{l}})\in V(\tau_{l+1}=\tau\times\{0\})$ and the $T$-action on the $1$-dimensional locus $V(\tau_{l+1})$ is trivial and connected. Indeed, the toric blow up $b_{\Sigma^{\prime}}$ restricts to ${\overline{V}(\tau_{l+1})}(\simeq\mathbb{A}^{1}_{\mathbbm{k}})\to p_{\sigma}\times\mathbb{A}^{1}_{\mathbbm{k}}(\subset{\rm Spec}(R)\times U_{\sigma})$ as an $T\times\mathbb{G}_{m}$-equivariant isomorphism, where $p_{\sigma}$ denotes the unique $T$-invariant closed point of $U_{\sigma}$, but the $T$-action on both the source and the target are trivial. Hence its base change gives an isomorphism $V(\tau_{l+1})\overset{\simeq}{\to}{\rm Spec}(R)\times p_{\sigma}\simeq{\rm Spec}(R)$ and the $T$-action on its source and target are again trivial. Hence, we claim that for such one dimensional $V(\tau_{l+1})$, we easily confirmed:
Claim 1.
$V(\tau_{l+1})$ is connected and $V(\tilde{\tau}^{\prime}_{l})$ is in the closure of $V(\tau_{l+1})$. The $T$-actions on it and the image of $M$ are both trivial.
Now we suppose $\tilde{f}(V(\tilde{\tau_{l}}))\in Z^{+}$ and obtain the contradiction as follows. Intuitively, the arguments is by “contagion of $Z^{+}$” to the image of the generic point of ${\rm Spec}(R)$. Firstly, if $\tilde{f}(V(\tilde{\tau_{l}}))\in Z^{+}$, it would imply
(9) $\tilde{f}(V(\tau_{l+1}))\subset Z^{+},$
from the above Claim 1. Furthermore, note that there is a natural $T$-equivariant morphism (retraction) $U_{\tau}\times{\rm Spec}(R)\to V(\tau_{l+1})\simeq{\rm Spec}(R)$. Hence, from the assumption of Definition 3.3 of higher $\Theta$-strata, that $\mathcal{Z}^{+}$ gives a union of connected components in ${\rm Map}(\Theta_{\tau},\mathcal{M})$, the connectedness of $U_{\tau}$ implies that $\tilde{f}(\eta_{R})\in Z^{+}$, where $\eta_{R}$ is the generic point of ${\rm Spec}(R)$. Hence, we get contradiction and it follows that $\tilde{f}(V(\tilde{\tau_{l}}))\notin Z^{+}$. From the proof, it also follows that $\tilde{f}(V(\tau_{l+1}))\cap Z^{+}=\emptyset$. We summarize here.
Claim 2.
$\tilde{f}(V(\tilde{\tau_{l}}))\notin Z^{+}$ and $\tilde{f}(V(\tau_{l+1}))\cap Z^{+}=\emptyset$.
On the other hand, note that $\tilde{f}(V(\tilde{\tau}^{\prime}_{0}))=f(\mathfrak{m})\in Z^{+}$ and that $\tilde{f}(V(\tau^{\prime}_{1}))\subset Z^{+}$. Therefore, there is some $i$ between them i.e., $i\in\{1,\cdots,l\}$ such that
  • $\tilde{f}(V(\tau^{\prime}_{i}))\subset Z^{+}$ and thus $\tilde{f}(V(\tilde{\tau}^{\prime}_{i}))\in Z^{+}$,
  • $\tilde{f}(V(\tilde{\tau}_{i-1}))\in Z^{+}$.
Suppose there is such $i$ as above with $i>1$. Note that the closure $\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}}$ is $V(\tilde{\tau}^{\prime}_{i})\cup V(\tau^{\prime}_{i})\cup V(\tilde{\tau}^{\prime}_{i-1})$. Hence, it follows that $\tilde{f}(\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}})\subset Z^{+}$.
Now we recall that $Z^{+}$ has the meaning of attracting locus towards $U_{\tau}$-direction, in the context of generalized versions of the Białynicki-Birula decomposition. The original statement in [BB73, 4.8] was for smooth $M$ (and $r=1$), but if $M$ is singular, it is later defined categorically ([HL14, §1.4, Proposition 1.4.1] for general $r$ and [Dri13, §0.2, §1.3, §1.4], [DG14, §1.4-§1.7]) as indeed it only maps to $M$ as an unramified monomorphism (cf. [HL14, Corollary 1.4.3, Remark 1.4.4], [DG14, 1.6.2, 1.7.4]) rather than an immersion. Nevertheless, from our assumption in Definition 3.3 that that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed immersion, its base change $Z^{+}\to M$ is also a closed immersion in our case.  [^8] In the context of algebraic reconstruction of K-moduli of Fano varieties, this issue is sorted out in [BHLLX21, Proposition 7.2] by using (the arguments of) [BLZ22, 5.2]. In particular, the action morphism
$\alpha\colon\mathbb{G}_{m}\times Z^{+}$ $\to M,$
$(t,x)$ $\mapsto t\cdot x$
extends to $Z^{+}\times\mathbb{A}^{1}_{t}\to M$. If we restrict to $t=0$ case and further to
$\tilde{f}(\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}})\times\{0\}\to M,$
its open locus $V(\tau^{\prime}_{i})$ maps to a point $\tilde{f}(V(\tilde{\tau}^{\prime}_{i}))$, hence $\alpha$ contracts $\tilde{f}(\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}})$ to a point in $M$.
Note that removal of $\tau^{\prime}_{i}$ from $\Sigma^{\prime}$ to a coarser fan, which we denote as $\Sigma^{\prime}(i)$, corresponds to a blow down of $\overline{\bar{V}(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}}$ in $T_{N}{\rm emb}(\Sigma^{\prime})$ and the above discusssion shows that $\mathcal{U}_{\Sigma^{\prime}}\to M$ descends to $\mathcal{U}_{\Sigma^{\prime}(i)}$. By repeating this, we obtain the following.
Claim 3.
There exists $\tilde{f}$ which satisfies
  1. ($\mathrm{i}$)
    $\tilde{f}(V(\tilde{\tau}^{\prime}_{i}))\not\subset Z^{+}$ for all $i=1,\cdots,l$,
  2. ($\mathrm{ii}$)
    but $\mathcal{U}_{\Sigma^{\prime}}\to M$ is still a morphism (not only a rational map).
Therefore, by setting $c_{1}$ as $c^{\prime}$ and $\tau^{\prime}_{1}$ as $\tau^{\prime}$, we complete the proof of Lemma 3.11. ∎
We continue the proof of Theorem 3.8. We write the simplicial subcone $\tau^{\prime}$ as $\sum_{i=1}^{r}\mathbb{R}_{\geq 0}v^{\prime}_{i}$, where $\{v^{\prime}_{i}\}_{i}\cup(\vec{0},1)$ are generators of the extremal rays of $\tau^{\prime}$. We denote $\tilde{\tau}_{1}$ (of the above proof of Lemma 3.11) as $\widetilde{\tau}^{\prime}$. Write $v^{\prime}_{i}=(v_{i},\frac{l_{i}}{m_{i}})$ with $l_{i},m_{i}\in\mathbb{Z}_{>0}$ and set $m:={\rm lcm}\{m_{i}\}$.
Take an extension of $R$ as follows. If $R$ is a discrete valuation ring, then consider $R\to R^{\prime}$ with ramification degree $m$, which is obtained as a localization of $R[x]/(x^{m}-\pi)$ at $(x)$. Note that $\{\frac{m_{i}v^{\prime}_{i}}{m}\}_{i}$s together with $(\vec{0},\frac{1}{m})$ gives a basis of a regular cone of $N\oplus\frac{1}{m}\mathbb{Z}$. From the construction, (base change of) $\tilde{f}$ is extendable to $\widetilde{\rm ST}_{R^{\prime}}(\tau^{\prime})\to M$ and ${\overline{\rm ST}}_{R^{\prime}}(\tau^{\prime})\to\mathcal{M}$ and the desired properties are satisfied from the construction and finish the proof of (i). Note that the extension $R\to R^{\prime}$ does not change when we shrink $\tau^{\prime}$, as it follows from the construction. If $R$ is not divisorial valution ring, then we instead apply Lemma 3.5.
Now we show the uniqueness assertion (iii). First, note that for each positive real number $c^{\prime\prime}$, if $(\xi,c^{\prime\prime})$ is contained in a rational hypersurface, then it is defined only by $c^{\prime\prime}$ because $\xi$ is assumed to be irrational if $r>1$. We recall that from the proof of Lemma 3.11, we have $\tilde{f}(\eta_{V(\tau^{\prime})})\notin Z^{+}$ and $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})\in Z^{+}$, where $p_{-\tau}$ denotes the $T$-invariant point of affine toric variety $U_{-\tau}$ over $R/\mathfrak{m}$. Here the meaning of $\lim$ is naturally justified by the extension of the $T$-action to the N-spectrum $\Delta_{-\xi}$ and leave the detailed definition to the readers. From these conditions, $c^{\prime}$ is determined canonically. Indeed, in case $c^{\prime\prime}>c^{\prime}$ and a rational polyhedral cone $\tau^{\prime\prime}$ which contains $(\xi,c^{\prime\prime})$, we would have $\tilde{f}(\eta_{V(\tau^{\prime\prime})})\in Z^{+}$ and in case $c^{\prime\prime}<c^{\prime}$ then $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})\notin Z^{+}$. Indeed, the latter limit is $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})=\tilde{f}(V(\hat{\tau}^{\prime\prime}))$, where $\hat{\tau}^{\prime\prime}$ is the cone of $\Sigma^{\prime}$ which contains $(\xi,(c^{\prime\prime},c^{\prime\prime}+\epsilon))$. However $\tilde{f}(V(\hat{\tau}^{\prime\prime}))$ is not contained in $Z^{+}$ from the higher $\Theta$-strata condition. Hence we complete the proof of the uniqueness assertion and hence Theorem 3.8 is now proved for the case of quotient stacks by algebraic split torus $T$.
Step 2.
In the second step, we consider the general quotient stack case i.e., when $\mathcal{M}$ is a quotient stack of the form $[M/G]$ where $M$ is an algebraic scheme over $\mathbbm{k}$ acted by an affine algebraic $\mathbbm{k}$-group $G$.
We essentially reduce the proof of the case to when $\mathcal{M}$ is a quotient stack of the form $M$ an algebraic scheme over $\mathbbm{k}$ acted by an algebraic $\mathbbm{k}$-split torus $T$ of dimension $r$.
Let us set $\mathcal{M}:=[M/G]$, where $M$ is a finite type $\mathbbm{k}$-scheme acted by an algebraic $\mathbbm{k}$-group $G$. Here we denote $m\colon T\times T\to T$ is the multiplication and $a\colon G\times M\to M$ is the action. We introduce the following concepts for the reduction to quotient stack by an algebraic split torus.
Property 1 (Local liftability).
  1. ($\mathrm{i}$)
    We call a higher $\Theta$-strata on $[Z^{+}/G]\subset\mathcal{M}$ for $G$-invariant closed subscheme $Z^{+}(\subset M)$ of type $\tau\subset N\otimes\mathbb{R}$ is (Zariski) locally liftable if the corresponding $\mathcal{Z}^{+}\subset\underline{\rm Map}(\Theta_{\tau}=[U_{\tau}/T],[M/G])$ with ${\rm ev}={\rm ev}_{(1,1,\cdots,1)}\colon\mathcal{Z}^{+}\simeq[Z^{+}/G]$ is given by the following data:
    • a disjoint union $M^{\prime}$ of Zariski open sets which covers $M$ (in particular, there is a surjective locally isomorphic morphism $c\colon M^{\prime}\twoheadrightarrow M$) with an action $a^{\prime}$ of $T$ on $M^{\prime}$ as $a^{\prime}\colon T\times M^{\prime}\to M^{\prime}$.
    • We have an algebraic $\mathbbm{k}$-morphism $M^{\prime}\to{\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ (homomorphism family), in other words $\psi\colon T\times M^{\prime}\to G$ which is homomorphism along each fiber of the second projection ${\rm pr}_{2}\colon T\times M^{\prime}\to M^{\prime}$ such that the following holds:
      1. (a)
        $c$ is $T$-equivariant with respect to $\psi$ in the sense that $a\circ c\circ{\rm pr}_{2}=c\circ a^{\prime}$ as morphisms from $T\times M^{\prime}$ to $M$.
      2. (b)
        $a^{\prime}|_{c^{-1}(Z^{+})}$ extends to $U_{\tau}\times c^{-1}(Z^{+})\to c^{-1}(Z^{+})$, which we still denote as $a^{\prime}$. Its composition with $c$ gives $\mathcal{Z}^{+}$ i.e., as a union of connected components of ${\rm Map}(\Theta_{\tau},\mathcal{M})$.
      Here, ${\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ naturally means the algebraic $\mathbbm{k}$-scheme representing homomorphisms of the algebraic $\mathbbm{k}$-groups.
  2. ($\mathrm{ii}$)
    We call a higher $\Theta$-strata on $[Z^{+}/G]\subset\mathcal{M}$ for $G$-invariant locally closed subscheme $Z^{+}(\subset M)$ of type $\tau\subset N\otimes\mathbb{R}$ is smooth (resp., fppf, étale) locally liftable if there is a smooth surjection (resp., faithfully flat surjection, étale surjection) $c\colon M^{\prime}\twoheadrightarrow M$ such that, the same holds as ($\mathrm{i}$) i.e., there exists $\psi$ which satisfies the above (a) and (b).
In practice, many examples of (higher) $\Theta$-strata satisfies the above Zariski local liftability condition ($\mathrm{i}$) and indeed we do not know any example which do not satisfies it. The main point of the above property is that it enables us the proof of Theorem 3.8 to the same arguments as Step 1 through the action $a^{\prime}$ on $M^{\prime}$. Note that this arguments do not depend on sophisticated results on algebraic stack theory, such as [AHHLR22], [AHR20, AHR19].
Now we prove the following.
Claim 4.
[…]
proof of Claim 4.
From the Białynicki-Birula decomposition [HL14, Theorem 1.4.7, 1.4.8], which extends [BB73, 4.8] (cf., e.g., also [Dri13, DG14]), the preimage of $\mathcal{Z}^{+}$ to the chart $X$, denoted as $Z^{+}(\subset X)$ satisfies that $Z^{+}/G\simeq\sqcup X^{\psi,+}/P_{\psi}$ for the attractor (blade in [HL14]) $X^{\psi,+}$ with respect to a homomorphism $\psi\colon T\to G$ in the notation of loc.cit. Note that due to the higher $\Theta$-strata condition that ${\rm ev}_{1}$ is closed immersion, for our proof, one can assume that $Z^{+}$ corresponds to only one $\psi$ as the representative of equivalence class modulo the Weyl type group $W$ of $G$. Then, from the same condition again, it follows that $Z^{+}=GX^{\psi,+}$ has a structure of closed subscheme of $X$ and there is an isomorphism $i\colon[Z^{+}/G]\simeq[X^{\psi,+}/P]$. Take the base change of $i$ by $(G\times X^{\psi,+})/P_{\psi}\to[X^{\psi,+}/P_{\psi}]$, we obtain an isomorphism  [^9] We thank T.Kinjo for the discussion on this. $(G\times X^{\psi,+})/P_{\psi}\simeq Z^{+}$. Note that $(G\times X^{\psi,+})/P_{\psi}$ (resp., $(G\times X)/P_{\psi}$ is a $X^{\psi,+}$-fiber bundle over $G/P$ (resp., $X$-fiber bundle) and in particular an algebraic space. By étale descent of the obvious fppf trivialization $(G\times X^{\psi,+})\to G\to G/P_{\psi}$ (resp., $(G\times X)\to G\to G/P_{\psi}$) of these fiber bundles, there is an étale surjection $B^{\prime}\twoheadrightarrow G/P$ so that the base change of $(G\times X)/P_{\psi}\to G/P_{\psi}$ (resp., $(G\times X^{\psi,+})/P_{\psi}\to G/P_{\psi}$ becomes trivial $X^{\psi,+}$-bundle i.e., isomorphic to product. That is,
(10) $X\times_{B^{\prime}}(G/P)$ $\simeq X\times B^{\prime},$
(11) $Z^{+}\times_{B^{\prime}}(G/P)$ $\simeq X^{\psi,+}\times B^{\prime}.$
By using $\psi\colon T\to G$, clearly there is an action on (10) which preserves its subset (11) and satisfies the desired Property 1 ($\mathrm{ii}$). ∎
Using the above Claim 4, now we can follow the arguments of Step 1. We take $M^{\prime}\twoheadrightarrow M$ as Claim 4 and consider a lift $p$ of $f(\mathfrak{m})$ for the closed point $\mathfrak{m}\in{\rm Spec}(R)$ as well as the morphism $f$. Note that since $f$ is formally étale and $R$ is complete, this is possible. Set $a^{\prime}(p_{\tau},p)=:p^{\prime}$. Then we consider the $T$-action $a^{\prime}$ on $M^{\prime}$ or a neighborhood of $p^{\prime}$ in particular, and resolve the indeterminancy of the rational map $U_{\tau}(\supset T)\times\Delta\dashrightarrow M^{\prime}$ by applying Lemma 3.11. We recall briefly what the claim of loc.cit implies: after replacing $\tau$ with a smaller enough subcone, it gives an appropriate rational polyhedral subcone of $\tau^{\prime}\subset\tau\times\mathbb{R}_{\geq 0}$ that projects isomorphically to $\tau^{\prime}$, set $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}$ and consider the corresponding affine toric variety $U_{\hat{\tau}^{\prime}}$, as well as its $1$-dimensional orbit $V(\tau^{\prime})(\simeq\mathbb{G}_{m})$ which corresponds to $\tau^{\prime}$. Then, $f$ extends to a certain $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$ for an appropriate $\hat{\tau}^{\prime}$ by Lemma 3.11, and by the exactly the same reasoning (Claim 3) one can take it so that $V(\tau^{\prime})\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$ maps outside $Z^{+}$ Note that this works by the strong condition Property 1 (b). In other words, by considering the action $a\circ(\psi\times{\rm id}_{M^{\prime}})\colon T\times M^{\prime}\to M$ and then apply Lemma 3.11 to it. Since $p^{\prime}\in M^{\prime}$, the existence of the extension of $f$ and the fact that $V(\tau^{\prime})\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$ maps outside $Z^{+}$ for appropriate $\hat{\tau}^{\prime}$ both directly follows.
Further, taking a rational ray $\mathbb{R}_{\geq 0}(v,d)\subset\tau^{\prime}$ for a primitive integral vector $v\in\tau$ and $d\in\mathbb{Z}_{>0}$, we obtain ${\rm Spec}\mathbbm{k}[[t^{\frac{1}{d}}]]\to U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$. Composing with the above $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$, we obtain the required replacement $f^{\prime}$ as a desired modification.
We write the following corollary to the above proof of Theorem 3.8, which is about how the semistable reduction of [AHLH23] changes along the variation of direction $\xi$.
For that, we prepare the following, to connect to the original $\Theta$-strata ([AHLH23]).
Definition 3.12 (Restriction to (usual $r=1$) $\Theta$-strata).
Consider $\xi\in\tau\subset N_{\mathbb{R}}$, following Notation 1, but we take $\xi$ to be integral i.e., $\xi\in N\cap\tau$ here. If we set $\iota\colon\mathbb{Z}\xi\hookrightarrow N$ the natural inclusion, note that there is a $({\rm exp}\otimes\iota)\colon\mathbb{C}\xi\to T(\mathbb{C})$ (compare Definition 2.4).
For a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$, consider restriction map $({\rm exp}\otimes\iota)^{*}\colon\mathcal{Z}\to{\rm Map}(\Theta_{\tau},\mathcal{M})$ and suppose it gives the $\Theta$-strata structure for each $\xi$ i.e., the images are connected. Then we call the obtained $\Theta$-strata as restriction along $\xi$ of the $\Theta$-strata $\mathcal{Z}$ and often denote as $\mathcal{Z}_{\xi}$.
Corollary 3.13 (of proof of 3.8 - effect of variation of $\xi$).
Fix a quotient stack $\mathcal{M}$ over $\mathbbm{k}$. We take a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}$ which has restrictions $\mathcal{Z}^{+}$ along each $\xi\in N\cap\tau$. As in Theorem 3.8 or [AHLH23, §6], take an arbitrary morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ where $R$ is a valuation $\mathbbm{k}$-algebra. Suppose that $f(\mathfrak{m})$ sits in a higher $\Theta$-stratum $\mathcal{Z}^{+}$ in $\mathcal{M}$ while $f(\eta_{R})$ is not in $\mathcal{Z}^{+}$.
Then, the limit of the obtained $\Theta$-semistable reduction $f|_{{\rm new},\xi}(0)$ by [AHLH23, Theorem 6.3] or Corollary 3.10 (and the one parameter family obtained $\mathcal{T}^{\prime\prime}$ in the proof of Theorem 3.8) varies only locally constantly along the change of $\xi$.  [^10] This phenomenon is later discussed in [BHLINK25] which appeared $8$ and a half months after our paper, for application to generalize Donaldson-Thomas type theory in a derived setup [BDINKP25]. Compare [BHLINK25, Constancy theorem, §6.1 and §7.3 case (i)].
It often occurs in various moduli problems that a priori we only have a uniform degeneration along $\xi$-direction but not along a higher dimensional toric variety $U_{\tau}$. Nevertheless, the following lemma builts the connection, following Proposition 2.2.
Lemma 3.14 (Limit mapping stack).
For an algebraic stack $\mathcal{M}$ over $\mathbbm{k}$ with affine diagonal, following Notation 1, take irrational $\xi\in N_{\mathbb{R}}$, set $\Gamma_{\geq 0}:=\langle M_{\geq 0},\xi\rangle\subset\mathbb{R}$, and consider the global N-spectrum $\Delta_{\xi}^{\rm gl}=\Delta_{\xi}^{\rm gl}(\mathbbm{k})$ as in Definition 2.1. Note that the algebraic split torus $T:=\mathbb{G}_{m}\otimes N$ naturally acts on $\Delta_{\xi}$.
Consider various rational polyhedral cone $\tau$ which contains $\xi$. The mapping stack from the quotient stack (a priori, fibered only over the étale site of $\mathbbm{k}$-schemes which are not necessarily finite type) $[\Delta_{\xi}^{\rm gl}/T]$ is a locally finite type algebraic stack over $\mathbbm{k}$ and has a description as follows.  [^11] See Notes added at the end of this paper, for the relation with newly appeared [BHLINK25].
(12) ${\rm Filt}_{\xi}(\mathcal{X}):$ $={\rm Map}(\Theta_{\xi}=[\Delta_{\xi}^{\rm gl}/T],\mathcal{X})$
[…]
Here, the injective system is given by the inclusion of the cones $\tau^{\prime}\supset\tau$ which induces the pull back morphism $\iota_{\tau,\tau^{\prime}}^{*}\colon{\rm Map}([\Theta_{\tau}^{\prime}/T],\mathcal{X})\to{\rm Map}([\Theta_{\tau}/T],\mathcal{X})$ through the morphism $\iota_{\tau,\tau^{\prime}}\colon U_{\tau}\to U_{\tau^{\prime}}$. The obtained restriction morphism $\iota_{\tau,\tau^{\prime}}^{*}$ is an open immersion of algebraic stacks so that we have the injective limit and the identification (13) makes sense.
In the sequel papers [Od24a, Od24c], we use the idea of the following proof efficiently rather than applying the statements.
Proof.
Note that ${\rm Filt}_{\xi}(\mathcal{X})$ is a priori only defined as a $2$-functor. Consider any base $\mathbbm{k}$-scheme $B^{\prime}$, on which $T$ acts, and take an object of the groupoid ${\rm Filt}_{\xi}(\mathcal{X})(B)$, and represent it as a $T$-equivariant morphism $\varphi\colon B\to\mathcal{X}$ after passing to an étale covering $B\to B^{\prime}$. We can and do assume that $B$ is an affine scheme $\mathop{\mathrm{Spec}}\nolimits(R)$ for a $\mathbbm{k}$-algebra $R$ and the map is represented by that of $B$ to a chart of $\mathcal{X}$, which we write $\mathop{\mathrm{Spec}}\nolimits(S)$, for simplicity. It corresponds to a ring homomorphism $S\to\mathbbm{k}[\Gamma_{\geq 0}]$. Now we exploit the similar discussions to Proposition 2.2. Since $S$ is finitely generated $\mathbbm{k}$-algebra, there is a finitely generated saturated submonoid $\Gamma^{\prime}_{\geq 0}\subset\Gamma_{\geq 0}$ so that the image of $S$ lies inside $\mathbbm{k}[\Gamma^{\prime}_{\geq 0}]$. This shows that
[…]
is surjective for each $k$-scheme $B$.
What remains is to show that $\iota_{\tau,\tau^{\prime}}^{*}\colon{\rm Map}([\Theta_{\tau}^{\prime}/T],\mathcal{X})\to{\rm Map}([\Theta_{\tau}/T],\mathcal{X})$ is an open immersion. For that take $T$-equivariant $B\times U_{\tau}\to\mathop{\mathrm{Spec}}\nolimits(S)$ for $B=\mathop{\mathrm{Spec}}\nolimits(R)$ and consider for which $\mathop{\mathrm{Spec}}\nolimits(\mathbbm{k}^{\prime})=b\in B$, it extends to $U_{\tau}^{\prime}\times b\to\mathop{\mathrm{Spec}}\nolimits(S)$. We consider the corresponding $\mathbbm{k}$-algebra homomorphism $f\colon S\to B[\tau^{\vee}\cap M]$. Take a finite generator $s_{1},\cdots,s_{m}$ of $S$ as $\mathbbm{k}$-algebra and consider all the coefficients $r_{i,\vec{m}}$ of $f(s_{i})$ of the monomial terms which correspond to $\vec{m}\in\Gamma^{\prime}\setminus\Gamma_{\geq 0}$. Then it is easy to see that $U_{\tau}\times b\to\mathop{\mathrm{Spec}}\nolimits(S)$ extends to $U_{\tau}^{\prime}\times b\to\mathop{\mathrm{Spec}}\nolimits(S)$ if and only if $r_{i,\vec{m}}(b)=0$ for all $i,\vec{m}$. Hence we complete the proof. ∎
An analytic version of our semistable reduction is as follows. Here we only discuss quotient stacks by algebraic tori, mainly due to the lack of enough references on analytic stacks (but cf., e.g., [Ino19a, Appendix A] and references therein) as it suffices for us for the moment.
Theorem 3.15 (Analytic analogue of Theorem 3.8 and Corollary 3.10).
Let $\Delta:=\{t\in\mathbb{C}\mid|t|<1\}$ and $M$ a complex variety acted algebraically by $T(\mathbb{C})^{\rm an}\simeq(\mathbb{C}^{*})^{r}$ i.e., it is the complex analytification of an algebraic $T$-variety.
Consider an arbitrary holomorphic map $\tilde{f}\colon\Delta\to M$ induces $f\colon\Delta\to\mathcal{M}:=[M/T(\mathbb{C})^{\rm an}]$ whose target is regarded as a complex analytic stack. Suppose that for a (algebraic) higher $\Theta$-stratum $\mathcal{Z}^{+}$ of type $\tau$ with ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=Z^{+}\subset M$, $\tilde{f}^{-1}(Z^{+})\subset\Delta$ is set theoritically $\{0\}\subset\Delta$.
Then, the following claims hold.
  1. ($\mathrm{i}$)
    We can modify $f$ to avoid the higher $\Theta$-stratum in the following sense. If we fix an irrational $\xi\in\tau\setminus N_{\mathbb{Q}}$, then the map $\overline{f}$ can be taken canonically.
    • For some $m\in\mathbb{Z}_{>0}$, we replace $\Delta$ by $\Delta^{\prime}:=\{s\in\mathbb{C}\mid|s|<1\}$ with $\Delta^{\prime}\twoheadrightarrow\Delta$ which maps $s$ to $t=s^{m}$,
    • replace $\tau$ by a smaller rational poyhedral cone of $N_{\mathbb{R}}$ (if necessary)
    • take a rational polyhedral cone $\tau^{\prime}$ of $N\oplus\mathbb{Z}$ which projects as PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$, as in Definition 3.1,
    so that the following hold:
    $\tilde{f}$ extends to a $T(\mathbb{C})^{\rm an}$-equivariant morphism ${\widetilde{\rm ST}}^{\rm an}_{\Delta^{\prime}}(\tau^{\prime})\to M$, hence extends $f$ to $\overline{f}\colon{\overline{\rm ST}}^{\rm an}_{\Delta^{\prime}}(\tau^{\prime})\to\mathcal{M}$, and maps $V(\tau^{\prime})$ outside $Z^{+}$ and there is a holomorphic map $\iota\colon\tilde{\mathbb{C}}^{*}\to M(\mathbb{C})$ which is equivariant with respect to the complex Lie group homomorphism $\tilde{R}_{-\xi}\colon\tilde{\mathbb{C}}^{*}\to T(\mathbb{C})$ (Definition 2.4) and satisfies that
    1. (A)
      $T(\mathbb{C})\cdot{\rm Im}(\iota)\supset\tilde{f}(V(\tau^{\prime})),$ which is disjoint from $Z^{+}$,
    2. (B)
      there is a sequence $t_{i}(i=1,2,\cdots)$ in $\mathbb{R}$ with $t_{i}\to 0(i\to\infty)$ such that $\lim_{i\to\infty}\iota(t_{i})\in Z^{+}.$ (Actually the above limit is $\tilde{f}(p_{\hat{\tau}^{\prime}})$ automatically.)
    3. (C)
      Each point of ${\rm Im}(\iota)$ is obtained as $\lim_{i\to\infty}t(i)\tilde{f}(s_{i})$ where $t(i)\in T(\mathbb{C})$, $s_{i}(\in\Delta)$, $s_{i}\to 0$ for $i\to\infty$.
  2. ($\mathrm{ii}$)
    Such $\iota$ which satisfies ($\mathrm{i}$B), ($\mathrm{i}$C) is unique, up to a translation by the action of $T(\mathbb{C})$,
  3. ($\mathrm{iii}$)
    In particular, as the projective limit of $\overline{f}|_{V(\tau^{\prime})}$, we obtain a canonical map
    (14) $f_{\rm new}\colon\Delta_{\xi^{\prime}}^{\rm conv}(\mathbb{C})\to\mathcal{M}$
    for some $\xi^{\prime}=(\xi,c)$ with $c\in\mathbb{R}_{>0}$, which we denote by $f_{\rm new}$. Note $f_{\rm new}$ maps the closed point (origin) outside $Z^{+}$. This is the desired canonical modification of $f$.
Proof.
Since the proof is along the same line as that of Step 1 of Theorem 3.8, we only note the technical differences and the way to overcome them. The assumption that $T(\mathbb{C})\curvearrowright M$ is algebraic implies that $U_{\tau}\otimes\Delta\dashrightarrow M$ is meromorphic hence the $T$-invariant toroidal resolution of indeterminancy exists similarly as Lemma 3.11. If we apply the same arguments as Theorem 3.8 for $R:=\mathbb{C}[[t]]^{\rm conv}=\mathcal{O}^{\rm an}_{\mathbb{C},0}$ (the convergent power series ring), the rest of the arguments is almost the same except for following points:
  1. (a)
    $\iota$ is obtained as a morphism from $\tilde{\mathbb{C}}^{*}$, rather than the N-spectrum algebraically.
  2. (b)
    the item ($\mathrm{i}$C) of Theorem 3.15 is slightly weaker than the algebraic version in Theorem 3.8 in that it only concerns a sequence of real numbers $\{t_{i}\}_{i}$. Thus the uniqueness statement ($\mathrm{ii}$) is a priori stronger.
Now we make comments to overcome these differences to complete the proof. Firstly, the difference (a) does not become a problem as we can simply take a (closed) point in $\tilde{f}(V(\hat{\tau}))$, consider the $T(\mathbb{C})$-action on it, take limit along $U_{-\tau}(\mathbb{C})$ to the $T(\mathbb{C})$-invariant point $p_{-\tau}$ and combine with $\tilde{R}_{-\xi}$.
For the difference (b) can be overcome as follows. The main points are that $\mathbb{C}^{*}$ acts continuously on $M(\mathbb{C})$ with respect to their complex analytic topolgies, so that any orbit of $U(1)\otimes N$ in $M(\mathbb{C})^{\rm an}$ is compact. First we take $c^{\prime}$ and $\tau^{\prime}$ as Lemma 3.11. From the first condition of item ($\mathrm{i}$B), there is some $c_{j}(j\in\{1,\cdots,l\})$ in the proof of Theorem 3.8 (before Lemma 3.11) so that the limit $\lim_{i\to\infty}\iota(t_{i})$ in the second item of ($\mathrm{i}$B) is $\tilde{f}(V(\sigma_{j}))$ for the $r+1$-dimensional cone $\sigma_{j}$ of $\Sigma^{\prime}$ which intersects with $(\xi,(c_{j},c_{j}+\epsilon))$ for $0<\epsilon\ll 1$.
In case $c_{j}>c^{\prime}$ and a rational polyhedral cone $\tau^{\prime}_{j}$ which contains $(\xi,c_{j})$, we would have $\tilde{f}(\eta_{V(\tau^{\prime}_{j})})\in Z^{+}$ and in case $c_{j}<c^{\prime}$ then $\lim_{t\in T(\mathbb{C}),t\to p_{-\tau}}t\cdot\tilde{f}(x)=\tilde{f}(V(\sigma_{j}))\notin Z^{+}$, for any $x\in V(\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))$. Since any orbit of $U(1)\otimes N$ in $M(\mathbb{C})^{\rm an}$ is compact, we have $\lim_{i\to\infty}\iota(t_{i})=\lim_{t\in T(\mathbb{C}),t\to p_{-\tau}}t\cdot\tilde{f}(x)=\tilde{f}(V(\sigma_{j}))$ for any $x\in V(\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))$. Therefore, our $c^{\prime}$ coincides with that of Lemma 3.11. Hence the rest of the proof works similarly as Theorem 3.8. ∎
Remark 3.16.
If we apply Theorem 3.8 with convergent N-ring $R=\mathbb{C}[[\Gamma_{\geq 0}]]^{\rm conv}$ of Definition 2.1 (ii), and use Lemma 2.5, we reprove Theorem 3.10 and even extends to general submonoid $\Gamma_{\geq 0}$. We wrote Theorem 3.8 as a special (simpler) case for simplicity.
3.3. Semistable reduction along higher $\Theta$-stratification: by repetition of Corollary 3.10
We use the following setup, which is a straightforward extension of the notion of $\Theta$-stratification of [HL14].
Definition 3.17 (cf., [HL14, §2.1] for $r=1$ case).
For an algebraic $\mathbbm{k}$-stack $\mathcal{M}$ with affine diagonal, consider the mapping stack ${\rm Map}(\Theta_{\tau},\mathcal{M})$ of $\Theta_{\tau}$ to $\mathcal{M}$ with various $N$ and $\tau$, as algebraic $\mathbbm{k}$-stacks, after [AHR20, 5.10, 5.11], [AHR19, 6.23]. Extending [HL14, Definition 2.1.2], a higher $\Theta$-stratification in $\mathcal{M}$ consists of
  1. ($\mathrm{i}$)
    a totally ordered 121212our order is converted from the notation in [HL14] for convenience to study problems related to K-stability, as in [BHLLX21] set $\Gamma$, which has a maximal element $1$,
  2. ($\mathrm{ii}$)
    closed131313it is a priori not necessarily reduced, although in our semistable reduction type theorem, the scheme structure is not really important substacks $\mathcal{M}_{\leq c}$ of $\mathcal{M}$ for $c\in\Gamma$ which satisfies that
    • they are monotonely increasing in the sense that we have $(\mathcal{M}_{\leq c})_{\rm red}\supset(\mathcal{M}_{\leq c^{\prime}})_{\rm red}$ if $c>c^{\prime}$ in $\Gamma$ (here, $(\mathcal{M}_{\leq c})_{\rm red}$ denotes the natural reduced subscheme structure on the same support),
    • $\bigcap_{c\in\Gamma}\mathcal{M}_{\geq c^{\prime}}=\emptyset$.
  3. ($\mathrm{iii}$)
    We write the open substack $\mathcal{M}_{\leq c}\setminus{\rm ev}_{1}(\mathcal{M}_{=c})$ as $\mathcal{M}_{<c}$, which we assume to coinside with $\bigcup_{c^{\prime}<c}\mathcal{M}_{\leq c^{\prime}}$, and we sometimes write ${\rm ev}_{1}(\mathcal{M}_{=c})\subset\mathcal{M}$ simply as $\mathcal{M}_{=c}$ if there is no possibility of confusion.
    Then, for each $c\in\Gamma$, there is a higher $\Theta$-strata structure of type $\tau$ (Definition 3.3) for some $N,\tau$, which we write as ${\rm ev}_{1}\colon\mathcal{M}_{=c}\hookrightarrow\mathcal{M}_{\geq c}$. Here, we do not require to identify $N,\tau$ for different $c$ and we allow any rational polyhedral $\tau$ for each $N,c$.
  4. ($\mathrm{iv}$)
    for any point $p\in|\mathcal{M}|$, $\{c\in\Gamma\mid p\notin|\mathcal{M}_{<c}|\}$ has a maximal element (the corresponding condition, for $r=1$ case, is assumed in [HL14] while it is called well-ordered in [BHLLX21].)
By repeating Theorem 3.8 and Corollary 3.10 for each higher $\Theta$-strata, we obtain their following version.
Theorem 3.18 (Semistable reduction along higher $\Theta$-stratification).
Following the notation of Definition 3.17, we take an algebraic stack $\mathcal{M}$ which is a quotient stack over $\mathbbm{k}$, and its higher $\Theta$-stratification $\{\mathcal{M}_{<c}\}_{c}$ as Definition 3.17.
Consider any $\mathbbm{k}$-morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ for a valuation $\mathbbm{k}$-ring $(R,\mathfrak{m})$ with its fraction field $K$, which satisfies that $f(\mathop{\mathrm{Spec}}\nolimits(K))\in\mathcal{M}_{=1}$ while $f(\mathfrak{m})\in\mathcal{M}_{<1}$. Then, there is a finite sequence of modifications of $f$ and $R$ which we denote as
$f_{i}\colon\mathop{\mathrm{Spec}}\nolimits(R_{i})\to\mathcal{M}(i=0,1,\cdots,l)$
with $R\to R_{i}$ such that
  • $R_{i}$ is a valuation ring with injective local morphisms $\mathbbm{k}\to R\to R_{i}$ (we denote the fraction field of $R_{i}$ as $K_{i}$ and the maximal ideal as $\mathfrak{m}_{i}$),
  • $f=f_{0}$,
  • $f|_{\mathop{\mathrm{Spec}}\nolimits(K_{i})}=f_{i}|_{\mathop{\mathrm{Spec}}\nolimits(K_{i})}$ for each $i$,
  • $f_{i}(\mathfrak{m}_{i})\in\mathcal{M}_{=c_{i}}$ with $c_{0}<c_{1}<\cdots<c_{l}=1$,
  • in particular, $f_{l}(\mathfrak{m}_{l})\in\mathcal{M}_{=1}$.
For each $f_{i}$, if we consider the source stack $\Theta_{i}\simeq\Theta^{r}$ and fix an isomorphism to $\Theta_{\tau_{i}}$ for some $N_{i},\tau_{i}$, together with choices $\xi_{i}\in\tau_{i}\setminus N_{\mathbb{Q}}$. Then the above sequence of modifications $f_{i}$ can be taken canonically.
Proof.
We apply Theorem 3.8 to $f$ with the higher $\Theta$-stratification on $\mathcal{M}_{={c_{0}}}$ and take the restriction of the obtained extension $\overline{f}$ to the closed substack corresponding to a ray $\mathbb{R}_{\geq 0}v$ in $v\in N_{\mathbb{Q}}\cap\tau^{\prime}$ in the proof of Theorem 3.8, and denote it by $f_{1}$. Then, we continue the process. Then, from the condition ($\mathrm{iv}$), it stops after finite times so that we are done. The last statement about the canonical choices of $f_{i}$ follows also from Theorem 3.8. We can characterize such $f_{i}$ by the natural condition in terms of $\xi_{i}$ by Theorem 3.8 but we avoid the lengthy repetition. ∎
Note that from the proof and canonicity statement of Theorem 3.8, $f_{i}$ are taken almost canonically in the sense that $f_{i}(\mathfrak{m})$ are unique, and have certain characterizations.
4. Applications
4.1. Properness of moduli of Calabi-Yau cones
Now we assume $\mathbbm{k}$ is algebraically closed field of characteristic 00 and see our first application of higher $\Theta$-semistable reduction (Theorem 3.8, Corollary 3.10, Theorem 3.18) on the moduli of Calabi-Yau cones. The Calabi-Yau cones, or equivalently K-polystable Fano cones, mean certain affine Calabi-Yau varieties with $\mathbb{R}_{>0}$-homogeneous metric and compatible algebraic torus action on them. As we discuss its moduli theorey thoroughly in [Od24a], we refer to op.cit for the details, and purpose here is to focus on the properness discussion and show that essentially only a part of Theorem 3.8 is enough for its establishment.
Let us clarify what it means and what we prove as a new claim.  [^14] This existed in appendix of [Od24c] v2 before, but we just moved it to here. To show its properness, as we discuss in §6 of [Od24a], the higher $\Theta$-strata appears. In this subsection, we prove that it is even Zariski locally liftable in the sense of Property 1 ($\mathrm{i}$). We follow the arguments and notation of op.cit, notably its §3.2 and §3.5.
Theorem 4.1 ([Od24a, 3.12]).
Fixing any positive integer $n$, $V$, Reeb vector field $\xi$, consider the moduli stack $\mathcal{M}_{V}^{\rm ss}(\xi)$ of $n$-dimensional K-semistable Fano cones of the minimized normalized volume $V$ and Reeb vector field $\xi\in N\otimes\mathbb{R}$. Then, the complement of $\mathcal{M}_{V}^{\rm ss}(\xi)\subset\mathcal{M}_{V}(\xi)$ admits even Zariski locally liftable (see Property 1 ($\mathrm{i}$)) higher $\Theta$-stratifiction so that $\mathcal{M}_{V}^{\rm ss}(\xi)$ is universally closed. Thus, $\mathcal{M}_{V}^{\rm ss}(\xi)$ satisfies the existence part of the valuative criterion of properness.
Proof.
We prove by the valuative criterion. Recall that $\mathcal{M}_{V}^{\rm ss}(\xi)$ is defined as a quotient stack $[B_{V}^{\rm ss}/G]$ in loc.cit, and $B_{V}^{\rm ss}$ can be enlarged to a larger $G$-acted algebraic $\mathbbm{k}$-scheme $B_{V}$ which is stratified as $B_{V}^{\rm ss}=B_{V}(0)\subset B_{V}(1)\subset\cdots\subset B_{V}(m^{\prime})=B_{V}$ by the value of the minimized normalized volumes again as in [Od24a, §3.2]. Thanks to the boundedness result [XZ24], now one can and do assume that each $B_{V}(i)$ contains all $n$-dimensional K-semistable Fano cone of the same volume density $V_{i}$, with $V=V_{0}>V_{1}>\cdots V_{m^{\prime}}$.
We take an arbitrary $\mathbbm{k}$-morphism ${\rm Spec}\mathbbm{k}[[t]]\to B_{V}$ such that the image $f(p)$ of the closed point $p$ lies in a connected component $Z^{+}$ of $B_{V}(i)\setminus B_{V}(i-1)$. (Recall that by the arguments in [Od24a], following [LX14], it follows that any $\mathbbm{k}$-morphism $\mathbbm{k}$-morphism ${\rm Spec}\mathbbm{k}((t))\to B_{V}^{\rm ss}$ can be extended to such for sufficiently large $B_{V}$ and $i$).
[…]
Then, if we restrict to $Z^{+}=B$, $\pi^{-1}(B)=:\mathcal{U}_{B}\to B$ has $(T\times)T_{B}$-action and canonical ideal sequences $I_{\lambda_{i}}\subset\mathcal{O}_{\mathcal{U}_{B}}$ for a discrete subset $\lambda_{i}\subset S(B)\subset\mathbb{R}_{\geq 0}$ with $0=\lambda_{0}<\lambda_{1}<\cdots$, which satisfies that $I_{\lambda_{i}}/I_{\lambda_{i+1}}$ is locally free $\mathcal{O}_{B}$-module of rank $r_{i}$ ([Che24, 1.1 (also §5)]). Here, $S(B)$ is the holomorphic spectrum (in the sense of [DS17]) corresponding to the $\pi$-fibers over $B$. Thanks to the fiberwise $T=N\otimes\mathbb{G}_{m}$-action, one can decompose $I_{\lambda_{i}}$ as $\oplus_{m\in M=N^{\vee}}I_{\lambda_{i},m}$ to the finite dimensional $T$-eigenspaces. Then, each $I_{\lambda_{i},m}/I_{\lambda_{i+1},m}$ is locally free $\mathcal{O}_{B}$-module of the finite rank $r_{i,m}$.
Now, consider the K-semistable degeneration of $[X]=f(p)\in B_{V}$ as $X\rightsquigarrow W$ and realize it as the $T_{B}$-action for some inclusion $T_{B}\subset G$ induced by the splitting of $\mathcal{O}_{X}\supset I_{\lambda_{1}}|_{X}\supset I_{\lambda_{2}}|_{X}\supset\cdots$. The limit $[W]$ still lies inside $B_{V}$ by the construction of $B_{V}$ mentioned above, as the minimized normalized volume remains the same as that of $X$ (cf., e.g., [XZ24]). We extend this inclusion $T_{B}\subset G$ as a homomorphism family locally around $[W]$ as follows. Since the $G$-action on $\mathcal{U}_{B}$ and the fiberwise $T$-action commutes, for each $\lambda_{i}$ and $m\in M^{\prime}$, we have a canonical $T_{B}(\times T)$-equivariant splitting of $\mathbbm{k}$-vector spaces $I_{\lambda_{i},m}|_{W}\twoheadrightarrow(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})|_{W}$ due to the complete reductivity of $T_{B}$. We extend it as follows. If we consider $I_{\lambda_{i},m}$ and $(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})$ as $\mathcal{O}_{B}$-modules, they are both locally free and inherit the $T_{B}$-action by the canonicity of $I_{\lambda_{i}}$s. Then, if we shrink $B$ to a small enough $T_{B}$-invariant open affine neighborhood of $[W]$ in $B$ by Sumihiro’s theorem, the $T_{B}$-equivariant vector bundles restricted to it are the direct sum of $T_{B}$-equivariant line bundles (cf., [IS15, 2.5], [Pay08, 2.1], [Kly89]).
By the same arguments as loc.cit, the $T_{B}$-equivariant surjection between $I_{\lambda_{i},m}\twoheadrightarrow(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})$ is $T_{B}$-equivariantly splittable by the complete reductivity of $T_{B}$ as $(I_{\lambda_{i},m}/I_{\lambda_{i+1},m})\to I_{\lambda_{i},m}$ which corresponds to $T_{B}$-semiinvariant fiberwise coordinates on $\mathcal{U}_{B}=\pi^{-1}(B)$. We extend it to $T_{B}$-semiinvariant functions on $\pi^{-1}(B_{V}^{o})$ for a $T_{B}$-invariant open affine neighborhood $B_{V}^{o}$ of $[W]$ in $B_{V}$ which vanish at the 00-section($\simeq B_{V}^{o}$). If we collect them for all $m\in M^{\prime}$ and many enough $i$, and compare with the embedding of $\mathcal{U}\to\mathbb{A}^{l}\times B_{V}$ (for the same $M^{\prime}$), then this gives family of linear transformation $g(b)\in G\subset{\rm GL}(l)$ for each $b\in B_{V}^{o}$, which a priori non-trivially varies over $B_{V}^{o}$.
Then the family of homomorphism $((Z^{+})^{\prime}:=)B_{V}^{o}\to{\rm Hom}(T_{B},G)$ given by $b\mapsto g(b)^{-1}T_{B}g(b)$ ensures the local liftability of the higher $\Theta$-strata structure on $Z^{+}=B$, so that Theorem 3.8 applies. By iterating the same procedure $i$-times (Theorem 3.18), we obtain the desired modification ${\rm Spec}\mathbbm{k}[[t]]\to B_{V}^{\rm ss}$ of $f$, so that we conclude the assertion by the valuative criterion. ∎
4.2. Properness of moduli of (singular) Kähler-Ricci solitons
In this subsection, we consider another application of our methods to the (expected) properness problem of moduli of compact shrinking Kähler-Ricci soliton (cf., [PSS23, GPSS18, CSW18, Ino22, HL20, BLXZ23]). As we mentioned in §2, this class of Kähler manifolds or their singular extensions, called compact shrinking ($\mathbb{Q}$-)Kähler-Ricci soliton is a certain class of log terminal Fano varieties, generalizes the class of K-polystable Fano varieties.
Here we discuss its moduli theory after the above references and we prove that expected properness is reduced to some natural but technical input from birational geometry of K-stability. This material is somewhat analogous to [Od24c, §2], and is even somewhat easier, since the objects are proper rather than affine. In this subsection, we assume the base field $\mathbbm{k}$ is algebraically closed field of characteristic 00.
Recall that the celebrated work of Chen-Sun-Wang on complex Fano manifolds [CSW18] provides another $2$-step degeneration
$X\rightsquigarrow(T\curvearrowright X^{\prime},\xi^{\prime})\rightsquigarrow(T\curvearrowright X^{\prime\prime},\xi^{\prime\prime}),$
where $(T\curvearrowright X^{\prime},\xi^{\prime})$ and $(T\curvearrowright X^{\prime\prime},\xi^{\prime\prime})$ are (modified) K-semistable and K-polystable in the sense of [BN14] respectively, hence the latter admitting (weak) Kähler-Ricci solitons as the polarized limit space (Gromov-Hausdorff limit) of the normalized Kähler-Ricci flow. Later it is more algebraically reformulated by the work of [DS20, Ino22, Ino19a, HL20, BLXZ23]. In particular, Dervan-Szekelyhidi [DS20] introduces  [^15] Here, we use the opposite sign convention of [DS20] following [HL20, BLXZ23]. the $h$-invariant $h(X)\in\mathbb{R}_{\leq 0}$ for any $\mathbb{Q}$-Fano variety $X$ ([HL20] for the kawamata-log-terminal singular pairs case).
So, given our previous subsection of the moduli of possibly irregular Calabi-Yau cones, especially the proof of the properness by using the higher $\Theta$-semistable reduction Theorem 3.8, it now becomes natural to expect that an analogous method may work for the moduli of Fano varieties with (weak) shrinking Kähler-Ricci solitons.
As the setup, let $N$ again be a free abelian group of finite rank $r$ and we consider an algebraic $\mathbbm{k}$-torus $T=N\otimes\mathbb{G}_{m}$ which acts on log terminal $\mathbb{Q}$-Fano varieties $X$ of dimension $n$, together with $\xi\in N\otimes\mathbb{R}$ so that the triple $(T\curvearrowright X,\xi)$ is K-semistable and its anticanonical volume $(-K_{X})^{n}$ is $V$ (or consider its natural logarithmic generalization, for fixed coefficients set). We consider their moduli Artin stack $\mathcal{M}^{s}_{n,V,T,\xi}$, as an easy proof of its existence is included in our Theorem 4.3 for readers convenience, following e.g., [PSS23, Ino22, BLXZ23]. We also fix $n,V,h$ and consider the finite type moduli Artin stack $\mathcal{M}_{n,V,h}$ (resp., $\mathcal{M}_{n,V,h^{+}}$), after [BLXZ23, 6.1, 6.4], of $n$-dimensional $\mathbb{Q}$-Fano varieties $X$ with anticanonical volume $(-K_{X})^{n}=V$ and $h(X)=h$ (resp., $h(X)\geq h$).
The latter is (essentially) constructed in [BLXZ23, §6, Theorem 6.1 (and 6.4)], as op.cit Theorem 6.1 proves the existence of $\mathcal{M}^{\rm Fano}_{n,V,(1,)h^{+}}$ which is our $\mathcal{M}_{n,V,h^{+}}$, and it follows from their Theorem 6.4 that $\mathcal{M}_{n,V,h}$ is a locally closed substack of $\mathcal{M}_{n,V,h^{+}}$. Indeed, by the same theorem, it even follows that $\mathcal{M}_{n,V,h^{+}}$ has a finite stratification by of $\mathcal{M}_{n,V,h^{\prime}}$ for finite possibilities of $h^{\prime}(\geq h)$. Now we hope that the following analogue of [Che24, 1.1], [Od24a, Theorem 3.5] hold.
This is a stronger version of properness conjecture of K-moduli of (singular) Kähler-Ricci solitons, as we reveal the relation in Theorem 4.3 below.
Conjecture 4.2 (Higher $\Theta$-stratification by $h$-invariant of Fano varieties).
The $h$-invariants of $\mathbb{Q}$-Fano families and the corresponding first step degeneration of [CSW18, BLXZ23] (as reformulated in [HL20, §2], and Example 2.19, Corollary 2.17) satisfies the following general properties.
  1. ($\mathrm{i}$)
    Consider any finite type $\mathbbm{k}$-scheme $S$ and a $\mathbb{Q}$-Gorenstein 161616locally stable in the sense of [Kol23] family $\pi\colon\mathcal{X}\to S$, whose $h$-invariants and volumes are both constant i.e., $h(X_{s})$, $(-K_{X_{s}})^{n}$ does not depend on closed points $s\in S$, where $X_{s}:=\pi^{-1}(s)$. Then, there is a finite rank free abelian group $N$ and a rational polyhedral cone $\tau\subset N\otimes\mathbb{R}$ together with $\xi\in S$ such that there is a $\mathbb{Q}$-Gorenstein family $\tilde{\pi}\colon\tilde{\mathcal{X}}\to U_{\tau}\times S$ which restricts to $\pi$ (resp., the corresponding first step degeneration of $X_{s}$ [CSW18, BLXZ23]) on $(1,\cdots,1)\times S$ (resp., $\Delta_{\xi}\times s\to U_{\tau}\times s$, on each $s$). Here, $U_{\tau}$ means the affine toric variety corresponding to $N_{\mathbb{R}}\supset\tau$ and $\Delta_{\xi}$ is the N-spectrum in §2 (cf., also [HL20, §2]).
  2. ($\mathrm{ii}$)
    For a positive integer $n$ (the dimension), real numbers $V$ (the anticanonical volume, hence positive) and $h$ (lower bound for the $h$-invariant), the finite type moduli Artin stack of $\mathbb{Q}$-Fano varieties $X$ with $h(X)\geq h$ $\mathcal{M}_{n,V,h^{+}}$ (cf., [BLXZ23, 6.1]) and its natural finite stratification $\{\mathcal{M}_{n,V,h^{\prime}}\}_{h^{\prime}(\geq h)}$ by the values of $h(X)$ ([BLXZ23, Theorem 6.4]) admits the structure of a (étale locally liftable) higher $\Theta$-stratification in the sense of §3.3, by encoding the aforementioned first step degeneration of $\mathbb{Q}$-Fano varieties [CSW18, BLXZ23].
Note that (the special case when $S$ is a smooth curve of) the above ($\mathrm{i}$) easily implies ($\mathrm{ii}$). Indeed, because the $H^{\rm NA}$-invariant for a special $\mathbb{R}$-test configuration only depends on the central fiber with the vector field (cf., [HL20, Example 2.32 (87), §2.5 (73)]), together with the fact that the Duistermaat-Heckman measure of $\mathbb{R}$-filtration is determined by the dimensions of the each filters (cf., [BC11], §2.6), such $\tilde{\pi}$ can exists only if $h(X_{s})$ is constant. The rest follows from the similar arguments as [Od24a, 3.13 (§3.3, §3.4)] after [HL14, AHR20]. See also a related criterion of the $\Theta$-stratification in [HL14], given by two conditions (S) and (B) therein.
We also note that the same arguments as the previous subsection or [BHLLX21, Conjecture 1.1 and around] (cf., also [LXZ22] whose main theorem is its solution) shows that the former ($\mathrm{i}$) follows from a certain finite generation of a graded ring whose ${\rm Proj}$ will give $\tilde{\pi}$ and its closed fibers, and hence it is again a problem of birational geometry and we expect it will be similarly solvable by using [BCHM10] and later developments.
Now discuss the item ($\mathrm{ii}$). If it holds, our higher $\Theta$-semistable reduction theorem 3.8 applies and we readily obtain its application as below. This is again parallel to the discussions in the previous subsection for the Calabi-Yau cone case and [Od24c] (see next subsection).
Theorem 4.3 (Properness of moduli of Kähler-Ricci soliton modulo finite generation problem).
Under the above setup, we assume the conjecture 4.2 holds. Then, both the moduli algebraic $\mathbbm{k}$-stacks $\mathcal{M}_{n,V,h}$ and $\mathcal{M}^{s}_{n,V,T,\xi}$ are universally closed. 171717i.e., satisties the existence part of the valuative criterion of properness of an algebraic stack. The actual properness requires separatedness in some of the literature.
Also, the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ of (modified) K-semistable pair $(X\curvearrowleft T,\xi)$ exists. If $\mathbbm{k}=\mathbb{C}$ and the moduli stack $\mathcal{M}^{s}_{n,V,T,\xi}$ admits a categorical good  [^18] We refer to [Stacksproject, 106.12] (resp., [Alp13]) for the definition of categorical moduli (resp., good moduli). Before, the author has often used the term “coarse moduli” for the exactly same meaning as the categorical moduli in older papers, i.e., in a broader sense than e.g., [Mum65], not assuming the condition of bijectivity of geometric points as in op.cit. moduli space $M^{s}_{n,V,T,\xi}$, its complex analytification $M^{s}_{n,V,T,\xi}(\mathbb{C})$ is compact.
Remark 4.4.
In particular, by setting $h:=0$, we give a yet another proof of the properness of K-moduli stack of $\mathbb{Q}$-Fano varieties (modulo Conjecture 4.2), different from [BHLLX21, LXZ22] which uses the $\delta$-invariant [FO18, BlJ20].
Proof.
Consider any stalk of the structure sheaf on a proper $\mathbbm{k}$-curve at a closed point and denote the DVR as $R$, its fraction field (the function field) as $K$. Take an arbitrary $\mathbbm{k}$-morphism $\varphi^{o}\colon{\rm Spec}K\to\mathcal{M}_{n,V,h_{1}=h}$, corresponding to a $\mathbb{Q}$-Fano family $\pi^{o}\colon\mathcal{X}^{o}\to{\rm Spec}(K)$. By [LX18], after a finite base change (we keep the notation of $R,K$ though), $\pi^{o}$ extends to a $\mathbb{Q}$-Fano family $\pi\colon\mathcal{X}\to{\rm Spec}(R)$. We suppose the $h$-invariant of the central fiber (resp., generic fiber) of $\pi$ is $h_{0}$ (resp., $h_{1}$). By [BLXZ23], we have $h_{0}\leq h_{1}$. Correspondingly, so that $\varphi^{o}$ extends to a $\mathbbm{k}$-morphism from ${\rm Spec}(R)$ to $\mathcal{M}_{n,V,h_{0}^{+}}$ for some $h_{0}\leq h_{1}$.
What we need to prove is that, if $h_{0}<h_{1}$, we have another extension of $\varphi^{o}$ to $\mathcal{M}_{n,V,h_{1}=h}$ as $\varphi\colon{\rm Spec}(R)\to\mathcal{M}_{n,V,h_{1}=h}$ possibly after finite extension of $R$. If conjecture (4.2) ($\mathrm{ii}$) is true, then the assertion is a direct corollary to the the higher $\Theta$-semistable reduction along the higher $\Theta$-stratification §3.3, Theorem 3.18. Note that by boundedness results obtained in [BLXZ23, Jia20], possible values of $h(X)$ for $X\in\mathcal{M}_{n,V,h^{+}}$ are finite.
The existence of $\mathcal{M}^{s}_{n,V,T,\xi}$ as an algebraic $\mathbbm{k}$-stack follows similarly as [Od24a]: it is equivalent to the stack of $T$-equivariant $\mathbb{Q}$-Gorenstein families of $n$-dimensional $\mathbb{Q}$-Fano varieties of volume $V$, whose $H^{\rm NA}$-minimizer is $\xi$, by its uniqueness ([HL20, 1.2], [BLXZ23, 1.2]), and the constructible lower semicontinuity of the $h$-invariant ([BLXZ23, 6.4]). Indeed, the whole stack $\mathcal{M}^{T}$ of $T$-equivariant $\mathbb{Q}$-Gorenstein families of $n$-dimensional $\mathbb{Q}$-Fano varieties $X$ of volume $V$ is an algebraic $\mathbbm{k}$-stack. This comes from the fact that the classical affine cone $C(-NK_{X}):={\rm Spec}\oplus_{m\geq 0}H^{0}(-mNK_{X})$ for a uniform sufficiently divisible positive integer $N$, together with the action of $T\times\mathbb{G}_{m}$ on it, recovers $X$. Indeed, the data $C(-NK_{X})\curvearrowleft(T\times\mathbb{G}_{m})$ forms an algebraic stack due to the existence of the (quasi-projective) multi-Hilbert scheme [HM04] and the hulls and husks construction by Kollár [Kol08]. Then our $\mathcal{M}^{s}_{n,V,T,\xi}$ is straightforwardly seen to be an open substack of $\mathcal{M}^{T}$ by the aforementioned results in [HL20, BLXZ23].
Take a connected component $\mathcal{M}$ of $\mathcal{M}^{s}_{n,V,T,\xi}$, put the corresponding $h$-invariant as $h$ and consider the connected component of $\mathcal{M}_{n,V,h}$ which includes the natural image of $\mathcal{M}$, which we denote as $\mathcal{M}_{h}$. Therefore, the forgetful morphism $\mathcal{M}\to\mathcal{M}_{h}$ between the two algebraic $\mathbbm{k}$-stacks is radiciel and finite. Since the latter is universally closed, so is the former. We complete the proof. ∎
By the above properness theorem, we can expect very similar methods of constructing the good moduli space $\mathcal{M}^{s}_{n,V,T,\xi}$ of (weak) Kähler-Ricci solitons, to the construction of the good moduli space of the moduli stack of the Calabi-Yau cones in [Od24a]. Also note that the above algebraic compactness result is compatible with the precompactness type result by [PSS23, GPSS18].
4.3. Bubbling Kähler metrics
In [Od24c], which appears a few weeks later than this paper v1, we provide an application of Theorem 3.10 which we briefly review. In particular, it constructs finite certain affine Calabi-Yau varieties from any (strict) degeneration of log terminal singularities. This conjecturally corresponds to the differential geometric bubbles for the degeneration of local Kähler-Einstein metrics (to local singular Kähler-Einstein current) by their metric rescale up. In that case, the algebraic stack we use is a quotient stack of the form $[M/T]$ with an algebraic torus $T$, hence our step 1 of Theoerem 3.10 is enough. One of the first examples in the $2$-dimensional case, for instance, include (underlying affine varieties of) ALE Kähler manifolds by Kronheimer ([Kro89a]). We hope to explore this bubbling phenomenon more thoroughly in the near future. See [Od24c] for more details. As its byproduct, for instance, we also prove a weaker version of resolution of (log terminal) singularities, or rather an alteration in a more canonical manner (Corollary 2.9) which hopefully shows its potential depth.
Acknowledgement
The author thanks E.Inoue and M.Jonsson for the comments on the manuscript. During the research, the author was partially supported by Grant-in-Aid for Scientific Research (B) 23H01069, 21H00973, Grant-in-Aid for Scientific Research (A) 21H04429, 20H00112, and Fund for the Promotion of Joint International Research (Fostering Joint International Research) 23KK0249. Part of this work is done during the author’s stay at Institut de mathématiques de Jussieu - Paris Rive Gauche. We appreciate the kind hospitality.
Notes added in arXiv v3: about [BHLINK25]
About $8$ and a half months after the first apperance on arXiv and submission of this paper, we learnt from the editor about preparation of [BHLINK25, §7.3] which somewhat overlaps our §3.2. We would like to express our gratitude to D. Halpern-Leistner and T. Kinjo for informing us of his work prior to its appearance on arXiv and, above all, for the invaluable and fruitful discussions on the possible connections and fascinating differences between our approaches. These discussions have been insightful, and we sincerely appreciate the time and effort. While there are notable overlaps of our papers in certain aspects, our methods and main applications in our minds are quite distinct at the same time. We hope that our works complement each other and contribute meaningfully to future research.
Below, we provide a list of comments on the differences and overlaps between our works, reflecting our discussions with Tasuki Kinjo and also Chenjing Bu, for which I am truly grateful.
  1. ($\mathrm{i}$)
    The “real $\Theta$-strata (or -stratification)” of [BHLINK25, §7.3.2, §7.3.3] corresponds to our higher $\Theta$-strata with the (possibly irrational) direction $\xi(\in\tau)$ as in our Theorem 3.8. Also, the usage of cones ($C$, $\sigma$ etc) in [BHLINK25] is in the same way as our $\tau$ so that their “cone filtration (op.cit §5.1)” is parallel to our higher $\Theta$-strata.
  2. ($\mathrm{ii}$)
    Hence, their [BHLINK25, Theorem 7.3.5] overlaps our Corollary 3.10 (also Theorem 3.8, Theorem 3.18) substantially and essentially gives a different proof. Despite the overlap, there are subtle differences even at the level of statements - our statements of Corollary 3.10 (and Theorem 3.18) and [BHLINK25, Theorem 7.3.5]. In our Corollary 3.10, unlike [BHLINK25, 7.3.5], we do not need the assumption on $\tau$ to be small enough (see the assumption on loc.cit (ii)). On our side, a certain smallness condition on the cone is only required for the existence statement of the extension $\overline{f}$ part (i.e., the stronger statements) in Theorem 3.8.
    In our semistable reduction statements in Corollary 3.10, a smaller cone does not appear and is not required.
  3. ($\mathrm{iii}$)
    Also, in our Theorem 3.8 and Corollary 3.10, we discuss families over more general valuation rings, which are not necessarily discrete valuation rings (DVR) while [AHLH23, §6] and [BHLINK25, §7.3] work over DVRs.
  4. ($\mathrm{iv}$)
    This seems to be the most crucial difference, at least from the perspective of the author. Our proofs and approaches to the aforementioned similar results - namely, our Corollary 3.10 (and Theorem 3.18) and [BHLINK25, Theorem 7.3.5] seem to be rather different.
    Indeed, we use more scheme theoretic techniques, notably Lemma 3.11, unlike [BHLINK25], we do not rely on [AHLH23]. Thus, we reprove [AHLH23, Theorem 6.3, 6.5] for quotient stack case.
    On the other hand, the proof of [BHLINK25, 7.3.5] uses more stack theoretic arguments and reduces to the original [AHLH23, Theorem 6.3, 6.5].
  5. ($\mathrm{v}$)
    More precisely, [BHLINK25] has two main theorems both under certain assumptions (cf., ($\mathrm{vii}$)) on the algebraic stacks - a certain local constancy of components of the filtration stack and the graded points stack in their §6.1 (Theorem 6.1.2), and finiteness of “special faces” in their §6.2 (Theorem 6.2.3, cf., also 6.2.5 and 6.2.8). Then, in the last subsection §7.3, they reduce the semistable reduction theorem similar to ours to [AHLH23, 6.3, 6.5] by applying those two theorems, in particular. Our Corollary 3.13 is also parallel to their two main theorems in a certain aspect.
  6. ($\mathrm{vi}$)
    Also, regarding our Lemma 3.14 (similar arguments are also done in [Od24a, §3.5]), the constancy theorem combined with the finiteness theorem imply that (13) in each (quasi-compact) component becomes actually stationary.
  7. ($\mathrm{vii}$)
    We assume $\mathcal{M}$ is global quotient stack in our theorems, while both [BHLINK25, 6.1.2 resp., 6.2.3] assumes a somewhat technical conditions on algebraic stacks that is to have “finite cotangent weights” (resp., to have only “quasi-compact graded points”). These are comparable to our work as quotient stacks are good examples which satisfy both conditions See op.cit §5.3.3 and 6.2.2 for the explanation and also [BHLINK25, Example §5.3.5] for non-quotient algebraic stack which do not satisfy their assumptions.
  8. ($\mathrm{viii}$)
    The discussions in [BHLINK25], notably its §5.2, partly relies on the theory of cotangent complex.